### v4

# Partial Law Invariance and Risk Measures

Abstract
We introduce the concept of partial law invariance, generalizing the concepts of law invariance and probabilistic sophistication widely used in decision theory, as well as statistical and financial applications. This new concept is motivated by practical considerations of decision making under uncertainty, thus connecting the literature on decision theory and that on financial risk management. We fully characterize partially law-invariant coherent risk measures via a novel representation formula. Strong partial law invariance is defined to bridge the gap between the above characterization and the classic representation formula of Kusuoka. We propose a few classes of new risk measures, including partially law-invariant versions of the Expected Shortfall and the entropic risk measures, and illustrate their applications in risk assessment under different types of uncertainty. We provide a tractable optimization formula for computing a class of partially law-invariant coherent risk measures and give a numerical example.
Keywords: Law invariance, probabilistic sophistication, ambiguity, coherent risk measures, source theory, Expected Shortfall
1 Introduction
Risk measures are popular tools in financial regulation and decision making. In finance, the usual interpretation of a risk measure is that it assigns a monetary amount to each financial position, representing the capital that should be held alongside the position to make it acceptable. In decision-theoretic terms, a risk measure is often interpreted as the unique certainty equivalent representing a preference relation over random variables, typically with some financially relevant properties. Risk measures such as the Value-at-Risk (VaR) and the Expected Shortfall (ES, also known as CVaR) are implemented in the financial regulatory framework of BCBS (2019), and they are also useful tools in optimization; see Föllmer and Schied (2016) and McNeil et al. (2015) for general treatments of risk measures in risk management and finance, and Shapiro et al. (2021) in the context of optimization.
Law invariance is the fundamental property that makes risk measures, as well as many decision models, applicable in practice. Once a reference probability model is specified, either through statistical estimation or expert opinion, law-invariant risk measures evaluate financial positions that share the same distribution under this probability model equally. This connects the theory of risk measures to statistics, thus allowing for a description of the probabilistic structure of financial positions, arguably a basic step in any quantitative financial analysis. In decision theory, law invariance is often known as probabilistic sophistication (Machina and Schmeidler (1992)). Although law invariance is a fundamental concept, it may not be universally desirable in all contexts. For instance, model misspecification, or model uncertainty, occurs when the assumed probabilistic model for various financial positions does not accurately represent the true, unknown, distribution of these positions. This can be problematic as the value of a law-invariant risk measure may differ significantly between the assumed model and the objective truth. For instance, during the 2008 financial crisis, the probability model used for default rates of mortgages was questionable, and the loss distribution for mortgage-backed securities was underestimated, leading to insufficient capital requirements for many companies to remain solvent (e.g., McNeil et al. (2015)). Although we take a risk measure perspective in this paper, the problem of model uncertainty extends beyond this and can affect any (usually statistical) process that relies on estimating a probabilistic model, followed by computing a quantity based on that model. We will generally refer to uncertainty in the probabilistic model as “model uncertainty”, which is also known as Knightian uncertainty or ambiguity in the literature. There are many avenues of research that account for model uncertainty in risk assessment in different ways. One area is distributionally robust optimization, in which optimization with risk measures is performed against a worst-case scenario; see e.g., Ghaoui et al. (2003), Natarajan et al. (2008), and Zhu and Fukushima (2009). Another area is to study risk measures formulated from multiple scenarios; see e.g., Delage and Li (2018), Delage et al. (2019) and Wang and Ziegel (2021). Model uncertainty is deeply rooted in decision theory (often called ambiguity there), with influential work pioneered by Schmeidler (1989) and Gilboa and Schmeidler (1989), followed by e.g., Klibanoff et al. (2005) and Maccheroni et al. (2006).
Most of the literature, including the works cited above, addresses model uncertainty through two approaches. The first approach begins with a reference probability model and a law-invariant risk measure. This risk measure would reflect an individual’s attitudes towards the financial positions if the individual had complete confidence in the reference probability model. Subsequently, to account for model uncertainty, the risk measure is aggregated over multiple alternative probability models, where each alternative model is motivated by the reference probability model. The aggregation is often, but not always, the worst-case aggregation, which is the approach in distributionally robust optimization, where the law-invariant risk measure is a commonly used one, such as VaR or ES. The second approach begins with a general risk measure that depends solely on the outcomes of the financial positions across possible future states of the world. Unlike the first approach, this approach does not assume a reference probability model, meaning the risk measure cannot be law invariant from the outset. This method circumvents model uncertainty by never supposing a model from the start, and it is consistent with the decision models of Schmeidler (1989) and Maccheroni et al. (2006). In the context of risk measures, this approach is quite common; see Föllmer and Schied (2016).
In this paper, we take a novel approach. We begin with a reference probability model but not a law-invariant risk measure. To account for model uncertainty, risk measures are assumed to satisfy law invariance only partially, in a sense to be specified later. The following example can motivate this idea. After estimating a probability model for various loss factors, it is natural that different factors will have different levels of model uncertainty. For example, this difference might result from a statistical inference procedure, where the confidence regions for parameters tied to certain loss factors could be larger or smaller than those for other loss factors, for instance, due to data scarcity. From this point, we can typically divide the loss factors into two categories: the loss factors for which we have no, or negligible, model uncertainty and those that have non-negligible model uncertainty. We will call all the events related to the loss factors in the first category the “model-certain source.” A financial position is said to depend on this source if its outcome is determined by events within this source. The next step is to select a risk measure for decision making. Requiring law invariance would not be able to distinguish factors in or outside the model-certain source, which is not desirable. However, for financial positions depending on the model-certain source, it is reasonable to measure their risk level solely based on distributions given by the probability model, as these positions involve no model uncertainty; thus, a partial form of law invariance is desirable. With this guiding principle in mind, this paper introduces a new property called partial law invariance. Fix a source, mathematically modeled by a sub-$\sigma$-algebra $\mathcal{G}$ of the whole model space. A risk measure is partially law invariant (with respect to $\mathcal{G}$) if it equates the risks of financial positions that (a) both belong to this source (i.e., $\mathcal{G}$-measurable) and (b) share identical distributions. This property reflects the discussion above with $\mathcal{G}$ being the model-certain source, but it is applicable beyond this interpretation. Although we take risk measures as the main objects of the paper, the concept of partial law invariance is naturally formulated for preference relations, applicable in broader contexts than financial risk assessment.
The approach of partial law invariance to modeling risk measures can address situations where some sources of randomness are more uncertain than others. Traditional approaches, such as distributionally robust optimization, cannot handle this setting, as these models of uncertainty cannot directly incorporate sub-$\sigma$-algebras.
The property of partial law invariance was inspired by, and closely related to, the recent paper of Baillon et al. (2023), where the authors formalize the idea of sources. In that paper, sources are defined as algebras of events, and the main idea is that uncertainty attitudes may differ between sources. We take a fixed source, corresponding to the source of risk in Baillon et al. (2023), allowing us to obtain mathematically tractable representation results. Our setting is later generalized to incorporate multiple sources and Section 4 discusses the connection between their setting and ours. A similar concept in decision theory involves having a $\sigma$-algebra with no ambiguity regarding the probability, as seen in the expected uncertainty utility model of Gul and Pesendorfer (2014).
We summarize the main contributions of the paper below. As the first contribution, we introduce the concept of partial law invariance as a new and relevant property for risk measures and preference relations in Section 2.1, with two substantive motivating examples in Sections 2.2 and 2.3, and a connection to stochastic dominance in Section 3.1. As the second contribution, a main theoretical result is presented, Theorem 1 in Section 3.2, which is a representation for coherent risk measures (Artzner et al. (1999)) satisfying partial law invariance. Third, generalizing the idea of a single source in Sections 23, the case of multiple sources is studied with a representation result Theorem 2 in Section 4, where we also explain the differences between the source theory of Baillon et al. (2023) and our approach. Fourth, we offer several extensions of the above results. A technical tool, called coherent adjustments, in Section 5.1 leads to an alternative representation of coherent risk measures in Theorem 3. This enables us to derive Theorem 4, which extends the well-known Kusuoka’s representation (Kusuoka (2001) and Frittelli and Rosazza Gianin (2005)) under a stronger notion of partial law invariance in Section 5.2. With no surprise to specialists in risk measures, most results can be smoothly generalized to the case of convex risk measures, and we present some of them in Section 5.3. Finally, formulas to compute the partially law-invariant ES are studied in Section 6, where we also present some applications and numerical examples. Section 7 concludes the paper. We will focus on an atomless probability space in the main part of the paper. The analysis in the case of a finite probability space is presented in the appendices, along with all proofs of technical results.
2 Partial law invariance: Definition and motivating examples
2.1 Risk measures, preference relations, and partial law invariance
We, throughout, fix a probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\mathbb{P}$ is our reference probability model. We denote by $L^{0}$ the set of measurable functions, where functions that agree $\mathbb{P}$-almost surely (a.s.) are identified. A generic set $\mathcal{X}\subseteq L^{0}$ is interpreted as the collection of financial positions of interest, and $X\in\mathcal{X}$ represents the loss of the financial position for various states of the world $\omega\in\Omega$ [^1] A positive value of $X$ represents a loss, and negative value of $X$ represents a profit; this is the standard loss/profit convention in the risk measure literature. It will always be assumed that $\mathcal{X}$ contains the constant random variables, identified with elements of $\mathbb{R}$. A risk measure is a mapping $\rho:\mathcal{X}\to\mathbb{R}$, with the main interpretation of the amount of capital needed to make a financial loss $X$ acceptable.
[…]
  • (C1)
    For all $X\in\mathcal{X}$, $\{c\in\mathbb{R}:X\succsim c\}$ and $\{c\in\mathbb{R}:c\succsim X\}$ are closed.
  • (C2)
    If $a,b\in\mathbb{R}$ and $a>b$, then $b\succ a.$ (Recall that elements of $\mathcal{X}$ represent losses.)
These assumptions allow us to guarantee the existence of a unique certainty equivalent, as described in the following standard result.
Proposition 1.
[…]
  1. $\mathcal{G}$-law invariance: $\rho(X)=\rho(Y)$ (resp. $X\sim Y$) for all $X,Y\in\mathcal{X}(\mathcal{G})$ that satisfy $X\buildrel\mathrm{d}\over{=}_{\mathbb{P}}Y$.
The property $\mathcal{F}$-law invariance is widely studied in the risk measure literature, called law invariance (or, probabilistic sophistication in decision theory). As far as we know, $\mathcal{G}$-law invariance with $\mathcal{G}\neq\mathcal{F}$ has not been studied in the literature. It is easy to see that law invariance is stronger than $\mathcal{G}$-law invariance. When making a contrast, we will refer to $\mathcal{G}$-law invariance as partial law invariance and to $\mathcal{F}$-law invariance as full law invariance, where the $\sigma$-algebras should be clear from the context. We also say that a set $\mathcal{X}\subseteq L^{0}$ is $\mathcal{G}$-law invariant if $X\in\mathcal{X}\cap L^{0}(\mathcal{G})$ and $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y\in L^{0}(\mathcal{G})$ imply $Y\in\mathcal{X}$.
To get a better intuition for partial law invariance, let us consider monotonicity of preferences and risk measures. For $X,Y\in\mathcal{X}$, we write $X\geq_{1}Y$ if $\mathbb{P}(X\leq x)\leq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}$, which is the usual first-order stochastic dominance (FSD). We can formulate two notions of monotonicity.
  1. Monotonicity: $X\succsim Y$ for all $X,Y\in\mathcal{X}$ with $Y\geq X$ (in the sense of $\mathbb{P}$-a.s. throughout).
  2. $\mathcal{G}$-FSD monotonicity: $X\succsim Y$ for all $X,Y\in\mathcal{X}(\mathcal{G})$ with $Y\geq_{1}X$.
The following simple result clarifies the connection between $\mathcal{G}$-FSD monotonicity and $\mathcal{G}$-law invariance.
Proposition 2.
[…]
Therefore, partial law invariance can be equivalently interpreted as consistency with respect to FSD over $\mathcal{G}$-measurable random variables.
The main motivations for partial law invariance, in the absence of full law invariance, are model uncertainty in risk evaluation, illustrated in Section 2.2, as well as ambiguity in decision making, illustrated in Section 2.3. In the following examples, and for the rest of the paper, we write $\sigma(\cdot)$ for the $\sigma$-algebra generated by either a random variable, a collection of random variables, or a collection of sets in $\mathcal{F}$.
2.2 Robust risk assessment under model uncertainty
Suppose a firm’s portfolio manager is tasked with making investment decisions involving the financial positions in $\mathcal{X}$. Among the random variables in $\mathcal{X}$, there is some financial position $Z$ where the manager knows that the distribution of $Z$ is accurate under the probability model, i.e., there is no (or negligible) model uncertainty; for instance, $Z$ may be the losses from a regular property insurance portfolio, where there is enough data to confidently estimate its distribution. The distribution of the other random variables $X$ under the probability model may be subject to a considerable amount of model uncertainty, such as the losses of a stock without a sufficiently long history of data. The manager needs to decide on a risk measure. In this context, law invariance naturally applies when dealing with positions $g(Z)$ that are functions of $Z$ because the manager is confident about their distributions. However, for random variables $X$ that are not $\sigma(Z)$-measurable, if $X\buildrel\mathrm{d}\over{=}_{\mathbb{P}}g(Z)$, it is unreasonable to insist that $X$ and $g(Z)$ are equally risky due to the model uncertainty associated with $X$; the same can be said for two general random variables $X$ and $Y$. In other words, it is natural to require $\sigma(Z)$-law invariance but not (full) law invariance.
Denote by $\mathcal{M}_{1}$ the set of probability measures on $(\Omega,\mathcal{F})$ that are absolutely continuous with respect to $\mathbb{P}.$ Due to the presence of model uncertainty and to be robust to model misspecification, the manager may choose a collection of probability measures $\mathcal{Q}\subseteq\mathcal{M}_{1}$ with the property that for all $\mu\in\mathcal{Q}$, $F_{\mu,Z}=F_{\mathbb{P},Z}$, where $Z$ is the loss whose distribution the manager is certain about, and $F_{\mu,Z}$ is the cumulative distribution function (cdf) of $Z$ under $\mu$. For example, the manager could take
$\mathcal{Q}=\left\{\mu\in\mathcal{M}_{1}\mid F_{\mu,Z}=F_{\mathbb{P},Z}\text{ and }d(\mu,\mathbb{P})<\epsilon\right\},$ (1)
where $\epsilon>0$ and $d:\mathcal{M}_{1}\times\mathcal{M}_{1}\to[0,\infty]$is some measure of divergence, such as a Wasserstein distance or the Kullback–Leibler divergence. For $X\in\mathcal{X}$, the manager needs to evaluate the risk of $X$ under each $\mu\in\mathcal{Q}$, and then aggregate these evaluations in some way. For instance, if the manager uses the mean to quantify risk and the worst-case approach of Gilboa and Schmeidler (1989) to aggregate the risk evaluations, then the resulting risk measure is $\rho:X\mapsto\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X]$. We can see that $\rho$ is $\sigma(Z)$-law invariant because $\rho(X)=\mathbb{E}^{\mathbb{P}}[X]$ for $X$ that is $\sigma(Z)$-measurable.
Using the mean as a risk measure is overly simplistic. A risk measure common in financial practice is the Expected Shortfall (ES also known as CVaR). ES serves as the standard risk measure for market risk in the banking regulatory framework of BCBS (2019); see Rockafellar and Uryasev (2002) for its optimization and Wang and Zitikis (2021) for an axiomatization. Denote by $L^{\infty}$ the set of essentially bounded random variables on $(\Omega,\mathcal{F},\mathbb{P})$. For a given probability measure $\mu\in\mathcal{M}_{1}$, the ES at level $\alpha\in[0,1)$ is given by
$\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left\{x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}[(X-x)^{+}]\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
The function $\mathrm{ES}^{\mu}_{\alpha}(X)$ can be defined as the average of the left-quantile function of $X$ under $\mu$ from $\alpha$ to $1$, and also on larger spaces in $L^{0}$; see e.g., Rockafellar and Uryasev (2002) and McNeil et al. (2015, Chapter 8). For simplicity, we write $\mathrm{ES}_{\alpha}=\mathrm{ES}_{\alpha}^{\mathbb{P}}$. Following the guidelines of BCBS (2019), for each $\mu\in\mathcal{Q}$, suppose that the decision maker uses $\mathrm{ES}_{\alpha}^{\mu}$ at level $\alpha=0.975$ to evaluate the risk of $X\in L^{\infty}$ under $\mu$. The resulting risk measure is the worst-case ES (see Zhu and Fukushima (2009)), given by
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}_{\alpha}^{\mu}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$ (2)
The risk measure $\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}$ with $\mathcal{Q}$ in (1) is $\sigma(Z)$-law invariant. Therefore, considering robust risk optimization will naturally lead to partial law invariance.
Another common way to assess model uncertainty is through the use of multiple prior probability measures. Law-invariant risk measures under multiple probability measures are studied by Wang and Ziegel (2021) under the name of scenario-based risk measures. It turns out that partial law invariance is closely related to scenario-based risk measures. Take a set $\mathcal{Q}\subseteq\mathcal{M}_{1}$. Wang and Ziegel (2021) defined the following property.
  1. $\mathcal{Q}$-basedness: $\rho(X)=\rho(Y)$ for all $X,Y\in L^{\infty}$ satisfying $F_{\mu,X}=F_{\mu,Y}$ for all $\mu\in\mathcal{Q}$.
This property means that if two risks are identically distributed under all probability measures of interest, then they are assessed as equally risky.
It is evident that law invariance is a particular case of both $\mathcal{G}$-law invariance, by choosing $\mathcal{G}=\mathcal{F}$, and $\mathcal{Q}$-basedness, by choosing $\mathcal{Q}=\{\mathbb{P}\}$. Moreover, law invariance is stronger in general than both $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness (if $\mathbb{P}\in\mathcal{Q}$). What is not so obvious is how $\mathcal{G}$-law invariance and $\mathcal{Q}$-basedness are connected, and we will first address this question. We denote by $L^{1}$ the set of integrable random variables on $(\Omega,\mathcal{F},\mathbb{P})$. For $\mu\in\mathcal{M}_{1}$, let $D_{\mu}=\mathrm{d}\mu/\mathrm{d}\mathbb{P}\in L^{1}$, that is, its Radon–Nikodym derivative with respect to $\mathbb{P}$.
Proposition 3.
If $\rho:L^{\infty}\to\mathbb{R}$ is $\mathcal{Q}$-based, then it is $\mathcal{G}$-law invariant for any $\mathcal{G}$ independent of $\sigma\left(D_{\mu}:\mu\in\mathcal{Q}\right)$.
This connection between $\mathcal{Q}$-basedness and $\mathcal{G}$-law invariance reinforces the idea that partial law invariance of a risk measure can serve as an indicator of sources without model uncertainty. In this context, observe that $\mathcal{Q}$ can be viewed as a set of competing probability models. The independence between $D_{\mu}$ and $\mathcal{G}$ for all $\mu\in\mathcal{Q}$ implies $\mu|_{\mathcal{G}}=\mathbb{P}|_{\mathcal{G}}$ for each $\mu\in\mathcal{Q}$. This means that all the competing probabilities in $\mathcal{Q}$ agree on $\mathcal{G}$, which can be directly interpreted as an absence of model uncertainty on $\mathcal{G}$, as in our motivating example.
2.3 Partial law invariance in decision models
[…]
We will now turn to some classical models in decision theory under ambiguity, such as those of Gilboa and Schmeidler (1989), Klibanoff et al. (2005) and Maccheroni et al. (2006). These models also naturally give rise to partially law-invariant functionals and preferences. We first describe a general framework that allows these models to be translated into our setting. Let $(\Omega_{0},\mathcal{F}_{0},\mathbb{P}_{0})$ denote the probability space representing the potential future states of the world. Suppose a decision-maker does not possess knowledge of $\mathbb{P}_{0}$ but has access to a lottery-generating device, modeled by $((0,1],\mathcal{B}((0,1]),\lambda)$, where $\lambda$ denotes the Lebesgue measure. Putting these two sources of uncertainty together, one arrives at the probability space
$(\Omega,\mathcal{F},\mathbb{P})=(\Omega_{0}\times(0,1],\mathcal{F}_{0}\otimes\mathcal{B}((0,1]),\mathbb{P}_{0}\times\lambda),$
[…]
$\Delta=\{\mu\in\mathcal{M}_{1}:\mu(A\times B)=\mu(A\times(0,1])\,\lambda(B)\text{ for all }A\in\mathcal{F}_{0}\text{ and }B\in\mathcal{B}((0,1])\},$
which represent the potential probability measures the decision maker needs to consider; note that the second marginal is always $\lambda$ since there is no ambiguity about the lotteries. Many examples in decision theory can be formulated via the above setting, and we list three below. In each model, $\ell:\mathbb{R}\to\mathbb{R}$ is a loss function, with the corresponding function $u(x)=-\ell(-x)$ being a von Neumann-Morgenstern utility function. (a) The maxmin expected utility of Gilboa and Schmeidler (1989), given by
$\rho(X)=\max_{\mu\in\mathcal{S}}\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu=-\min_{\mu\in\mathcal{S}}\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu,$
where $\mathcal{S}\subseteq\Delta$; (b) The smooth ambiguity model of Klibanoff et al. (2005), given by
$\rho(X)=\int_{\Delta}\phi\left(\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu\right)Q(\mathrm{d}\mu)=-\int_{\Delta}\psi\left(\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu\right)Q(\mathrm{d}\mu),$
where $\phi$ is an increasing transformation, $\psi(x)=-\phi(-x)$, and $Q$ is a probability measure over $\Delta$; (c) The variational preferences model of Maccheroni et al. (2006), given by
$\rho(X)=\max_{\mu\in\Delta}\left(\int_{\Omega}\ell(X)\leavevmode\nobreak\ \mathrm{d}\mu-c(\mu)\right)=-\min_{\mu\in\Delta}\left(\int_{\Omega}u(-X)\leavevmode\nobreak\ \mathrm{d}\mu+c(\mu)\right),$
where $c:\Delta\to[0,\infty]$ satisfies $c(\mu)=0$ for some $\mu\in\Delta$. Because measures in $\Delta$ have the same second marginal $\lambda$, all of the above models satisfy $\rho(Z)=\mathbb{E}^{\mathbb{P}}[\ell(Z)]$ for each lottery $Z\in L^{\infty}(\mathcal{G})$, and hence they are $\mathcal{G}$-law invariant, but clearly not law invariant in general.
[…]
3 Coherent risk measures and associated preferences
We will study partial law invariance in the framework of coherent risk measures and their relationship to preferences. We focus on coherent risk measures for three reasons: First, they are prominent in financial regulation and portfolio optimization (e.g., McNeil et al. (2015) and Föllmer and Schied (2016)). Second, they give rise to tractable representation results that provide insights into the structure and implications of partial law invariance. Third, they can be translated into the max-min preferences of Gilboa and Schmeidler (1989) with suitable interpretation. For the rest of the paper, we will assume that $\mathcal{X}=L^{\infty}$, as is common in the risk measure literature.
3.1 Preferences and risk measures
A coherent risk measure (Artzner et al. (1999)) is a mapping $\rho:L^{\infty}\to\mathbb{R}$ satisfying the following properties.
  • Cash invariance: $\rho(X+c)=\rho(X)+c$ for all $X\in L^{\infty}$ and $c\in\mathbb{R}$;
  • Monotonicity: $\rho(X)\leq\rho(Y)$ for all $X,Y\in L^{\infty}$ with $X\leq Y$;
  • Convexity: $\rho(\lambda X+(1-\lambda)Y)\leq\lambda\rho(X)+(1-\lambda)\rho(Y)$ for all $X,Y\in L^{\infty}$ and $\lambda\in[0,1]$;
  • Positive homogeneity: $\rho(\lambda X)=\lambda\rho(X)$ for all $X\in L^{\infty}$ and $\lambda\geq 0$.
Moreover, a convex risk measure is a mapping that satisfies the first three properties in the above list, and a monetary risk measure is a mapping that satisfies the first two properties in the above list; see Frittelli and Rosazza Gianin (2002) and Föllmer and Schied (2002). For background on risk measures, see Appendix A.
A common continuity assumption in the risk measure literature is Fatou continuity. For $\rho:L^{\infty}\to\mathbb{R}$, we say that $\rho$ is Fatou (continuous) if for all uniformly bounded sequences $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ that converge $\mathbb{P}$-a.s. to some $X\in L^{\infty}$, it holds that $\rho(X)\leq\liminf_{n\to\infty}\rho(X_{n}).$ Fatou continuity is essential because a risk measure $\rho:L^{\infty}\to\mathbb{R}$ is Fatou and coherent if and only if there exists $\mathcal{S}\subseteq\mathcal{M}_{1}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (3)
see Proposition 10 in Appendix A for a rigorous statement of this result. This representation is significant for two reasons. First, it links coherent risk measures to the maxmin expected utility model of Gilboa and Schmeidler (1989). With a given loss function $\ell$, the random variables $\ell(X)$ can be interpreted as a subjective loss; symmetrically, for a utility function $u$ and random wealth $X$, $u(X)$ is known as the util of $X$. Applying coherent risk measures to these subjective losses (or “dis-utils”) agrees with the maxmin expected utility model of Gilboa and Schmeidler (1989). Second, the worst-case expectation is evaluated with respect to a set of probability measures instead of finitely additive probabilities (which is the case without Fatou continuity). This distinction highlights the practical relevance of Fatou coherent risk measures, as statistical procedures yield probability measures rather than additive probabilities.
Preference relations that are represented by coherent risk measures will be called CRM preference relations, and they are described by a few properties; see Drapeau and Kupper (2013) and Proposition 12 in Appendix C. For CRM preference relations, $\mathcal{G}$-law invariance is connected to second-order stochastic dominance (SSD), which we show below. For $X,Y\in L^{\infty}$, we write $X\geq_{2}Y$ if $\mathbb{E}[\ell(X)]\geq\mathbb{E}[\ell(Y)]$ for all increasing convex functions $\ell:\mathbb{R}\to\mathbb{R}$. In the literature, the common definition is that for two random variables $Z$ and $W$ representing gains, $Z$ dominates $W$ in SSD if $\mathbb{E}[u(Z)]\geq\mathbb{E}[u(W)]$ for all increasing concave functions $u$. This is equivalent to $-W\geq_{2}-Z$, via the connection $\ell(x)=-u(-x)$, where $-W$ and $-Z$ represent losses. Next, we define SSD consistency.
  1. $\mathcal{G}$-SSD monotonicity: $X\succsim Y$ for all $X,Y\in L^{\infty}(\mathcal{G})$ with $Y\geq_{2}X$.
In general, $\mathcal{G}$-SSD monotonicity implies $\mathcal{G}$-FSD monotonicity because $\geq_{1}$ is stronger than $\geq_{2}$. These two notions of monotonicity are equivalent for CRM preferences, proved via standard arguments in the literature.
Proposition 4.
[…]
In the rest of the paper, we will focus on risk measures instead of preference relations. Given a monetary risk measure $\rho:L^{\infty}\to\mathbb{R}$, we call $\mathcal{A}_{\rho}=\{X\in L^{\infty}\mid\rho(X)\leq 0\}$ the acceptance set of $\rho$. Acceptance sets are a standard way to describe monetary risk measures; for some classical results regarding acceptance sets, see Appendix A. The following proposition relates partial law invariance and the acceptance set in the case of a monetary risk measure.
Proposition 5.
A monetary risk measure $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{A}_{\rho}$ is $\mathcal{G}$-law invariant.
3.2 Representation of partially law-invariant coherent risk measures
Our next task is to understand the structure of all $\mathcal{G}$-law invariant coherent risk measures. First, we need some notation. We denote by $\mathcal{M}(\mathcal{G})$ (resp. $\mathcal{M}_{1}(\mathcal{G})$) the set of finite signed measures (resp. probability measures) on $(\Omega,\mathcal{G})$ that are absolutely continuous with respect to $\mathbb{P}|_{\mathcal{G}}$. We write $\mathcal{M}=\mathcal{M}(\mathcal{F})$, and $\mathcal{M}_{1}=\mathcal{M}_{1}(\mathcal{F})$ for simplicity and to stay consistent with the notation from earlier. If $\mu\in\mathcal{M}_{1}(\mathcal{G})$, we will continue to use the notation $D_{\mu}$ to denote the Radon–Nikodym derivative with respect to $\mathbb{P}|_{\mathcal{G}}$, as $\mathrm{d}\mu/\mathrm{d}\mathbb{P}|_{\mathcal{G}}\in L^{1}(\mathcal{G})$. For $\mathcal{S}\subseteq\mathcal{M}(\mathcal{G})$ or $\mathcal{S}\subseteq\mathcal{M}$, we write $\widehat{\mathcal{S}}=\{D_{\mu}:\mu\in\mathcal{S}\}$, which is contained in $L^{1}(\mathcal{G})$ or $L^{1}$. Furthermore, in both cases, $\mathcal{S}$ is said to be $\mathcal{G}$-law invariant if the set $\widehat{\mathcal{S}}$ is $\mathcal{G}$-law invariant (defined in Section 2.1). For notational simplicity, when $\mu\in\mathcal{M}_{1}(\mathcal{G})$ we define $\mathbb{E}^{\mu}$ on $L^{\infty}$ by $\mathbb{E}^{\mu}[X]=\mathbb{E}[D_{\mu}X]$. For $\mathcal{S}\subseteq\mathcal{M}$, we write $\mathcal{S}^{\mathcal{G}}=\{\mu|_{\mathcal{G}}:\mu\in\mathcal{S}\}\subseteq\mathcal{M}(\mathcal{G})$. Similarly, for $\mathcal{B}\subseteq L^{1}$, we write $\mathcal{B}^{\mathcal{G}}=\{\mathbb{E}[X|\mathcal{G}]:X\in\mathcal{B}\}\subseteq L^{1}(\mathcal{G})$. There is no ambiguity when we use the notation ^$\widehat{\mathcal{R}}^{\mathcal{G}}$ because (^$(\widehat{\mathcal{R}})^{\mathcal{G}}=\widehat{(\mathcal{R}^{\mathcal{G}})}$. Given $\mu\in\mathcal{M}_{1}(\mathcal{G})$, let $\mathcal{E}(\mu)=\{\nu\in\mathcal{M}_{1}:\nu|_{\mathcal{G}}=\mu\}$. Clearly, $\mathcal{E}(\mu)$ depends on $\mathcal{G}$, which will be implicitly fixed in all places. If $\mu\in\mathcal{M}_{1}$, we will denote $\mathcal{E}(\mu|_{\mathcal{G}})$ by $\mathcal{E}(\mu)$ for simplicity.
Let $\mathfrak{X}$ and $\mathfrak{Y}$ be two vector spaces. Given a dual pairing $\left(\mathfrak{X},\mathfrak{Y},\langle\;,\;\rangle\right)$ let $w(\mathfrak{X},\mathfrak{Y})$ denote the weak topology on $\mathfrak{X}$. This makes $\left(\mathfrak{X},w(\mathfrak{X},\mathfrak{Y})\right)$ a locally convex (Hausdorff) topological vector space. Given some $\mathcal{R}\subseteq\mathfrak{X}$, let $\operatorname*{conv}(\mathcal{R})$ denote the convex hull of $\mathcal{R}$. It is well known that $\left(L^{\infty},\mathcal{M},\langle\;,\;\rangle\right)$ is a dual pairing under $\langle X,\mu\rangle=\mathbb{E}^{\mu}[X]=\mathbb{E}\left[D_{\mu}X\right].$ It is also well known that $\left(L^{\infty},L^{1},\langle\;,\;\rangle\right)$ is a dual pairing under $\langle X,Y\rangle=\mathbb{E}[XY].$ Indeed, $w\left(L^{\infty},\mathcal{M}\right)$ and $w\left(L^{\infty},L^{1}\right)$ coincide since $\left(\mathcal{M},w\left(\mathcal{M},L^{\infty}\right)\right)\cong\left(L^{1},w\left(L^{1},L^{\infty}\right)\right)$ through the mapping $\mu\mapsto D_{\mu}$.
For a Fatou coherent risk measure $\rho:L^{\infty}\to\mathbb{R}$, its supporting set is defined as:
$\mathcal{S}_{\rho}=\left\{\mu\in\mathcal{M}_{1}:\mathbb{E}^{\mu}[X]\leq\rho(X)\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ X\in L^{\infty}\right\}.$
Using the subjective loss interpretation for the random variables, we see that the supporting set is the set of priors in the the maxmin expected utility model of Gilboa and Schmeidler (1989).
We are now ready to present a general representation for Fatou $\mathcal{G}$-law invariant coherent risk measures reminiscent of representation (3).
Theorem 1.
For a mapping $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
    The risk measure $\rho$ is Fatou continuous, $\mathcal{G}$-law-invariant, and coherent;
  2. (ii)
    there exists a convex $\mathcal{S}\subseteq\mathcal{M}_{1}$ such that the $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closure of $\mathcal{S}^{\mathcal{G}}$ is $\mathcal{G}$-law invariant and
    $\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (4)
  3. (iii)
    the $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closure of $(\mathcal{S}_{\rho})^{\mathcal{G}}$ is $\mathcal{G}$-law invariant.
We say that $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ is almost $\mathcal{G}$-law invariant if the $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closure of $\mathcal{S}$ is $\mathcal{G}$-law invariant as in parts (ii)–(iii) of Theorem 1. It is well-known that a Fatou coherent risk measure $\rho$ is law invariant if and only if its supporting set $\mathcal{S}_{\rho}\subseteq\mathcal{M}_{1}$ is law invariant; see Föllmer and Schied (2016, Theorem 4.59). Theorem 1 is reminiscent of this result, but with some subtle differences; in particular, one needs to take the restriction of elements of $\mathcal{S}$, as well as a closure of the resulting set. The first operation is clearly needed, and the second one seems essential. For a discussion of this condition and a proof of Theorem 1, see Section C.2.
Example 1.
Let $\rho$ be the mapping $X\mapsto\mathbb{E}^{\nu}[X]$ where $\nu\in\mathcal{E}(\mathbb{P})$. Clearly, $\rho$ is $\mathcal{G}$-law invariant, and the only possible set $\mathcal{S}$ in (4) is given by $\{\nu\}$. We have $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$, which is almost $\mathcal{G}$-law invariant.
Example 2.
Using the subjective loss interpretation for the random variables, we see the preference relations related to the maxmin expected utility model of Gilboa and Schmeidler (1989) are $\mathcal{G}$-law invariant if and only if $\mathcal{S}^{\mathcal{G}}$ is almost $\mathcal{G}$-law invariant, where $\mathcal{S}$ is the set of priors. In the specific example given in Section 2.3, it is straightforward to show that $\mathcal{S}^{\mathcal{G}}=\{\tilde{\lambda}\}$, where $\tilde{\lambda}(\Omega_{0}\times B)=\lambda(B)$ for all $B\in\mathcal{B}((0,1]).$ Thus $\mathcal{S}^{\mathcal{G}}$ is almost $\mathcal{G}$-law invariant.
Theorem 1 also gives a feasible way to construct $\mathcal{G}$-law-invariant risk measures. We can take any law-invariant coherent risk measure $\tilde{\rho}$ on $L^{\infty}(\mathcal{G})$ (e.g., $\mathrm{ES}_{\alpha}$), and for each $\mu\in\mathcal{S}_{\tilde{\rho}}\subseteq\mathcal{M}_{1}(\mathcal{G})$ we take any set $\mathcal{S}(\mu)\subseteq\mathcal{E}(\mu)$. Setting
$\rho(X)=\sup_{\nu\in\mathcal{S}}\mathbb{E}^{\nu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ where }\mathcal{S}=\bigcup_{\mu\in\mathcal{S}_{\tilde{\rho}}}\mathcal{S}(\mu),$ (5)
yields a $\mathcal{G}$-law-invariant risk measure. Moreover, $\rho$ coincides with $\tilde{\rho}$ on $L^{\infty}(\mathcal{G})$. For instance, the risk measure $\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}$ in Section 2.2 coincides with $\mathrm{ES}_{\alpha}$ on $L^{\infty}(\mathcal{G})$.
3.3 Two constructions
Partial law invariance of coherent risk measures naturally appear in several concrete settings, which we discuss below. The first class concerns those built on conditional expectations.
Let $\tau:L^{\infty}\to L^{\infty}(\mathcal{G})$ be a mapping such that $\tau(X)=X$ for all $X\in L^{\infty}(\mathcal{G})$; an example would be $\tau(X)=\mathbb{E}[X|\mathcal{G}]$ or, more generally, conditional convex risk measures as defined in Detlefsen and Scandolo (2005). Consider risk measures $\rho$ of the form
${\rho}(X)=\tilde{\rho}(\tau(X)),\hskip 3.0ptX\in L^{\infty},$ (6)
where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is a law-invariant mapping. Clearly, $\rho$ defined in (6) is $\mathcal{G}$-law invariant because $\rho(X)=\tilde{\rho}(\tau(X))=\tilde{\rho}(X)=\tilde{\rho}(Y)=\rho(Y)$ for any $X,Y\in L^{\infty}(\mathcal{G})$ with $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$.
Remark 1.
There is a large literature on conditional and dynamic risk measures, which also involve a fixed sub-$\sigma$-algebra $\mathcal{G}$; see Riedel (2004), Detlefsen and Scandolo (2005), and Filipović et al. (2012). A conditional risk measures is a mapping $\rho:L^{\infty}\to L^{\infty}(\mathcal{G})$ satisfying some specific properties. We emphasize that this paper deals with the static setting because our risk measures take real values. However, it is possible, as shown above, to use conditional coherent risk measures to construct $\mathcal{G}$-law-invariant coherent risk measures, but our mathematical results do not overlap with the results in the above literature. Moreover, the motivation for partial law invariance differs from that of dynamic risk measures, although both involve sub-$\sigma$-algebras.
Functionals of the form (6) can be found in the two-stage market evaluation of Pelsser and Stadje (2014) and the market-consistent valuation of Dhaene et al. (2017). Setting $\tau(X)=\mathbb{E}[X|\mathcal{G}]$, equation (6) offers a straightforward method for constructing partially law-invariant risk measures. Specifically,
$\rho(X)=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}]),\quad X\in L^{\infty}.$ (7)
Coherent risk measures of the form (7) are characterized by the following result.
Proposition 6.
For a Fatou coherent risk measure $\rho:L^{\infty}\to\mathbb{R}$, equivalent are:
  1. (i)
    There exists a Fatou coherent risk measure $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ such that (7) holds;
  2. (ii)
    there exists $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ such that (7) holds;
  3. (iii)
    $\rho(X)=\rho(\mathbb{E}[X|\mathcal{G}])$ for all $X\in L^{\infty}$;
  4. (iv)
    $\widehat{\mathcal{S}}_{\rho}\subseteq L^{1}(\mathcal{G})$.
Moreover, if any of the above conditions are satisfied, $\rho$ is $\mathcal{G}$-law invariant if and only if $\tilde{\rho}$ is law invariant.
The risk measure in (7) further satisfies a stronger form of partial law invariance, which will be studied in Section 5.2.
Example 3.
The idea of using conditional expectations can be generalized further to incorporate different sub-$\sigma$-fields (see Section 4). Let $\mathcal{H}$ be a sub-$\sigma$-algebra independent of $\mathcal{G}$. Let $\rho_{1}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ and $\rho_{2}:L^{\infty}(\mathcal{H})\to\mathbb{R}$ both be coherent risk measures, with the former being law invariant. Define the following risk measure
$\rho(X)=\rho_{1}(\mathbb{E}[X|\mathcal{G}])\vee\rho_{2}(\mathbb{E}[X|\mathcal{H}]),\hskip 3.0ptX\in L^{\infty},$
where $a\vee b$ denotes the maximum between two real numbers $a$ and $b$. To show that $\rho$ above is $\mathcal{G}$-law invariant, we first notice that for $X$ that is $\mathcal{G}$-measurable, we have $\rho(X)=\rho_{1}(X)\vee\mathbb{E}[X].$ Coherence and law invariance of $\rho_{1}$ implies $\rho_{1}(X)\geq\mathbb{E}[X]$ for $X\in L^{\infty}(\mathcal{G})$; for this result, see Theorem 4.3 of Bäuerle and Müller (2006). Therefore $\rho(X)=\rho_{1}(X)$ for all $X\in L^{\infty}(\mathcal{G})$, which implies that $\rho$ is $\mathcal{G}$-law invariant.
The second class of partially law-invariant coherent risk measures arises from statistical functions and tail risk measures, similar to the example of robust ES in Section 2.2.
Let $\mathcal{P}_{c}(\mathbb{R})$ denote the set of Borel probability measures on $\mathbb{R}$ with compact support. A statistical function is mapping $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$. We say that a statistical function is coherent if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is coherent, where $X_{\#}\mathbb{P}\in\mathcal{P}_{c}(\mathbb{R})$ is the distribution of $X$ under $\mathbb{P}$. We say that a statistical function is Fatou if the functional on $L^{\infty}$ given by $X\mapsto\gamma(X_{\#}\mathbb{P})$ is Fatou.
Proposition 7.
Given a statistical function $\gamma$ and $\mathcal{S}\subseteq\mathcal{E}(\mathbb{P})$, the risk measure $\rho:L^{\infty}\to\mathbb{R}$ given by
$\rho(X)=\sup_{\mu\in\mathcal{S}}\gamma(X_{\#}\mu),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}$ (8)
is Fatou, $\mathcal{G}$-law invariant, and coherent.
Using Proposition 7, we can establish a connection between a “bottom-up” approach to addressing model uncertainty and the current framework. Suppose that an agent begins with a probability space $(\Omega,\mathcal{G},\mathbb{P}_{0})$, where they are certain about their probabilistic model $\mathbb{P}_{0}$ on $\mathcal{G}$. The agent wishes to extend this model to a finer-$\sigma$-algebra $\mathcal{F}\supseteq\mathcal{G}$, on which they have model uncertainty. To start, she defines a reference probability $\mathbb{P}$ on $\mathcal{F}$, such that $\mathbb{P}|_{\mathcal{G}}=\mathbb{P}_{0}$, to identify null sets. From here, the agent chooses a set of competing probabilistic models $\mathcal{S}\subseteq\mathcal{E}(\mathbb{P}_{0})$. For financial positions in $L^{\infty}(\mathcal{G})$, the agent intended to use the law-invariant risk measure $\rho:L^{\infty}(\mathcal{G})\to\mathbb{R}$ given by $\tilde{\rho}(X)=\gamma(X_{\#}\mathbb{P}_{0}),$ $X\in L^{\infty}(\mathcal{G})$, where $\gamma$ is a statistical function that is Fatou and coherent. The risk measure defined in equation (8) would be a natural risk measure to use on $L^{\infty}$. According to Proposition 7, this risk measure is Fatou, $\mathcal{G}$-law invariant, and coherent. For a similar approach in the dynamical setting, see Epstein and Schneider (2003) and Hansen and Sargent (2022).
Example 4.
For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\}$ and let $\tilde{\mathbb{E}}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ be the expectation statistical function (mapping a distribution to its mean). Define $\rho:L^{\infty}\to\mathbb{R}$ by
$\rho(X)=\sup_{\mu\in\mathcal{S}}\tilde{\mathbb{E}}(X_{\#}\mu)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
In this case, $\rho$ is of the form in (4) and (8). As $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$ is almost $\mathcal{G}$-law invariant, we can also verify that $\rho$ is $\mathcal{G}$-law invariant from Theorem 1. Take any $\sigma$-algebra $\mathcal{H}$ independent of $\mathcal{G}$, such that $(\Omega,\mathcal{H},\mathbb{P})$ is atomless (assumed to exist; this corresponds to a notion of conditional atomless property; see Remark 4 in Appendix B). We claim
$\rho(X)=\mathbb{E}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho(X)=\mathrm{ES}_{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ Y\in L^{\infty}(\mathcal{H}).$ (9)
Furthermore, these are the extreme values for random variables with a particular distribution. That is, if $Q\in\mathcal{P}_{c}(\mathbb{R})$ and $X\in L^{\infty}$ with $X_{\#}\mathbb{P}=Q$, we claim
$\tilde{\mathbb{E}}(Q)\leq\rho(X)\leq\tilde{\mathrm{ES}}_{\beta}(Q),$
where $\tilde{\mathrm{ES}}_{\beta}:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ is the ES statistical function. The $\sigma$-algebras $\mathcal{G}$ and $\mathcal{H}$ are extreme as $\rho(X)=\tilde{\mathbb{E}}(Q)$ if $X\in L^{\infty}(\mathcal{G})$ and $\rho(X)=\tilde{\mathrm{ES}}_{\beta}(Q)$ if $X\in L^{\infty}(\mathcal{H})$. For proofs of the above claims, see Section C.3.
By allowing $\mathcal{S}$ to vary but maintaining the condition that $\mathcal{S}^{\mathcal{G}}=\{\mathbb{P}|_{\mathcal{G}}\}$, we allow the risk measure to behave “nicely” on $L^{\infty}(\mathcal{G})$, while leaving flexibility outside this $\sigma$-field. Certainly, it is not necessary to require $\mu\in\mathcal{E}(\mathbb{P})$ for all $\mu\in\mathcal{S}$ (like $\mathcal{S}$ in Example 4); according to Theorem 1, it is necessary and sufficient to take $\mathcal{S}^{\mathcal{G}}$ to be almost $\mathcal{G}$-law invariant. Indeed, generalizing Example 4, we can have $\rho$ behave like a law-invariant risk measure $\widehat{\rho}$ on $L^{\infty}(\mathcal{G})$ and like a tail risk measure generated by $\widehat{\rho}$ on $L^{\infty}(\mathcal{H})$ in the sense of Liu and Wang (2021), this is illustrated by Example 5.
Example 5.
Given a Fatou law-invariant coherent risk measure $\widehat{\rho}:L^{\infty}\to\mathbb{R}$, there exists a Fatou coherent statistical function $\gamma:\mathcal{P}_{c}(\mathbb{R})\to\mathbb{R}$ such that $\widehat{\rho}(X)=\gamma(X_{\#}\mathbb{P})$ for all $X\in L^{\infty}$. For $\beta\in[0,1)$, define $\mathcal{S}=\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\}$. By Proposition 7, the risk measure $\rho$ given by (8) is Fatou, $\mathcal{G}$-law invariant, and coherent. We claim that the risk measure $\rho$ satisfies
$\rho(X)=\widehat{\rho}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \rho(X)=\widehat{\rho}^{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{H}),$
where $\widehat{\rho}^{\beta}$ is the $\beta$-tail risk measure generated by $\widehat{\rho}$. See Section C.3 for a proof of this claim.
4 Multiple sources
We first clarify the connection between partial law invariance and the source theory of Baillon et al. (2023). In the source theory, the decision maker is assumed to have a specific form of preference relation, and on the source $\mathcal{G}$ of risk, [^2] Baillon et al. (2023) used algebras instead of $\sigma$-algebras, which is a technical point and does not affect intuition. the decision maker is law invariant with respect to an objective probability. Our concept of partial law invariance assumes a probability $\mathbb{P}$ on the entire $\sigma$-algebra $\mathcal{F}$, which is interpreted as objective on $\mathcal{G}$. The requirement on $\mathcal{G}$ for the preference relation is the same in both frameworks.
The two frameworks differ on how random variables outside $\mathcal{G}$ are treated. In our framework, losses outside $L^{\infty}(\mathcal{G})$ are connected to those in $L^{\infty}(\mathcal{G})$ via financial properties like monotonicity and convexity, but we do not directly specify how their distributions are used in evaluating the risk. In the source theory, multiple sources are specified, and the decision maker is law invariant with respect to a (possibly different) subjective probability on each of them. [^3] In source theory, the natural property of monotonicity (i.e., $X\succsim Y$ if $X\leq Y$ pointwise, representing losses) across different sources (i.e., $X$ and $Y$ may not be measurable to the same source) seems to be difficult to embed.
Thus far, we have only considered one source (sub-$\sigma$-algebra) on which the risk measure is law invariant. In practical applications, there may be multiple sources of randomness on which the risk measure is law invariant. Extending our framework to multiple sources, we consider a similar idea as in Baillon et al. (2023), to let the decision maker be law invariant on each of the sources, but with respect to the common probability $\mathbb{P}$. This leads to the following formal property. Let $\Sigma(\mathcal{F})$ denote the set of all atomless sub-$\sigma$-algebras of $\mathcal{F}$ and $\mathfrak{G}\subseteq\Sigma(\mathcal{F})$. For a mapping $\rho:L^{\infty}\to\mathbb{R}$, we say that $\rho$ is $\mathfrak{G}$-law invariant if it is $\mathcal{G}$-law invariant for each $\mathcal{G}\in\mathfrak{G}$. When $\mathfrak{G}=\{\mathcal{G}\}$, $\mathfrak{G}$-law invariance is the same as $\mathcal{G}$-law invariance.
A natural corollary of Theorem 1 characterizes all Fatou $\mathfrak{G}$-law invariant coherent risk measures.
Theorem 2.
Let $\mathfrak{G}\subseteq\Sigma(\mathcal{F})$. A mapping $\rho:L^{\infty}\to\mathbb{R}$ is a Fatou $\mathfrak{G}$-law-invariant coherent risk measure if and only if
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}$
holds for some convex $\mathcal{S}\subseteq\mathcal{M}_{1}$ such that the $\mathcal{S}^{\mathcal{G}}$ is almost $\mathcal{G}$-law invariant for each $\mathcal{G}\in\mathfrak{G}$.
Appendix D contains an example of a $\mathfrak{G}$-law-invariant coherent risk measure given by $\rho(X)=\sup_{\mathcal{G}\in\mathfrak{G}}\rho_{\mathcal{G}}(\mathbb{E}[X|\mathcal{G}])$, where each $\rho_{\mathcal{G}}$ is law invariant and coherent, along with its representation in the form of Theorem 2.
There are at least two stronger yet natural ways of specifying partial law invariance across multiple sources, which we discuss below.
  1. (a)
    We may further require that the same risk evaluation applies across different $\mathcal{G}\in\mathfrak{G}$, leading to the condition that $\rho(X)=\rho(Y)$ for all $X,Y\in\bigcup_{\mathcal{G}\in\mathfrak{G}}L^{\infty}(\mathcal{G})$ that satisfy $X\buildrel\mathrm{d}\over{=}Y$. Note that $\bigcup_{\mathcal{G}\in\mathfrak{G}}L^{\infty}(\mathcal{G})$ is not necessarily a space or even a convex set, making the property potentially cumbersome to work with.
  2. (b)
    To obtain a proper linear space, we can consider the sub-$\sigma$-algebra $\sigma_{\mathfrak{G}}=\sigma\left(\bigcup_{\mathcal{G}\in\mathfrak{G}}\mathcal{G}\right)$, and work with $\sigma_{\mathfrak{G}}$-law invariance.
It is clear that both (a) and (b) are stronger than $\mathfrak{G}$-law invariance, and moreover, (b) is stronger than (a). In Example 6 below, we illustrate a situation in which (b) is not desirable, whereas (a) may be reasonable.
Example 6.
In the context of robust risk aggregation (see e.g., Embrechts et al. (2013, 2015), Bernard et al. (2014) and the references therein), the marginal distributions of $d$ random losses $X_{1},\dots,X_{d}$ are known, but their joint distribution is subject to substantial model uncertainty. In this case, it is natural to assume partial law invariance for $\mathfrak{G}=\{\sigma(X_{i}):i=1,\dots,d\}$, but not on the larger $\sigma$-algebra $\sigma(X_{1},\dots,X_{d})$, because the distributions of $X_{1}$ and $X_{2}$ are certain, but the distribution of $X_{1}+X_{2}$ is not. A concrete model is described below. Let $(\Omega,\mathcal{F})=\left(\mathbb{R}^{d},\mathcal{B}\left(\mathbb{R}^{d}\right)\right)$ and $\mathbb{P}$ be the multivariate Gaussian measure with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$. Let $\tilde{\Sigma}\in\mathbb{R}^{d\times d}$ be given by $\tilde{\Sigma}_{ij}=\mathds{1}_{\{i=j\}}\sigma_{i}\sigma_{j}$. For $\lambda\in[0,1]$, let $\mu_{\lambda}\in\mathcal{M}_{1}$ be the multivariate Gaussian measure with mean $m$ and covariance matrix $\lambda\Sigma+(1-\lambda)\tilde{\Sigma}$. Let $X_{1},\dots,X_{d}$ be the random variables given by the projections and $\mathfrak{G}=\left\{\sigma(X_{i}):i=1,\dots,d\right\}.$ For $\alpha\in[0,1)$, define the risk measure
$\rho(X)=\sup_{\lambda\in[0,1]}\mathrm{ES}^{\mu_{\lambda}}_{\alpha}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
For $i\in\{1,...,d\}$, let $\nu_{i}$ be the Gaussian measure on $\mathbb{R}$ with mean $m_{i}$ and variance $\sigma_{i}^{2}$. For all $\lambda\in[0,1]$ and $B\in\mathcal{B}(\mathbb{R})$, $\mu_{\lambda}\left(X_{i}^{-1}(B)\right)=\nu_{i}(B)$. Hence, $\rho$ is a $\mathfrak{G}$-law invariant coherent risk measure, and further it satisfies property (a) but not property (b).
Our current formulation of $\mathfrak{G}$-law invariance has an advantage over property (a), that is, the flexibility for modeling the situation where different sources are associated with different levels of model uncertainty (ambiguity), illustrated by Example 7 below.
Example 7.
Consider $\mathfrak{G}=\{\mathcal{G}_{0},\mathcal{G}_{1},\dots,\mathcal{G}_{n}\}$ with their pairwise intersection being generated by constants (an example would be that they are independent). Our interpretation is that $\mathcal{G}_{0}$ has no model uncertainty and each $\mathcal{G}_{i}$ is associated with different levels of model uncertainty. Suppose that the decision maker uses the multiplier preferences of Hansen and Sargent (2001) on each $L^{\infty}(\mathcal{G}_{i})$, $i\in\{0,1,\dots,n\}$, represented by
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}\left\{\mathbb{E}^{\mu}[\ell(X)]-\frac{1}{\theta_{i}}R(\mu,\mathbb{P})\right\}=\frac{1}{\theta_{i}}\log\mathbb{E}\left[e^{\theta_{i}\ell(X)}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G}_{i}),$ (10)
where $\theta_{i}\geq 0$, $\ell$ is an increasing loss function and $R(\mu,\mathbb{P})$ is the relative entropy of $\mu$ with respect to $\mathbb{P}$. The second equality in (10) is known as the Donsker–Varadhan variational formula; see e.g., Strzalecki (2011) for this formulation of the multiplier preferences via $\ell(x)=-u(-x)$ as in Section 2.3. When $\theta_{i}=0$, $\rho(X)=\mathbb{E}[\ell(X)]$, which corresponds to the limit as $\theta_{i}\downarrow 0$. The formulation (10) is well defined because the intersections of $L^{\infty}(\mathcal{G}_{i})$ and $L^{\infty}(\mathcal{G}_{j})$ for $i\neq j$ only contain constants $c\in\mathbb{R}$, for which $\rho(c)=\ell(c)$. In this model, $\theta_{i}$ represents the magnitude of model uncertainty in $\mathcal{G}_{i}$, with a larger value representing more severe model uncertainty; note that $\rho$ in (10) is increasing in $\theta_{i}$. We may assume $\theta_{0}=0$ because $\mathcal{G}_{0}$ has no model uncertainty (but it is not necessary for the discussions here). For $X$ outside $\bigcup_{i=0}^{n}L^{\infty}(\mathcal{G}_{i})$, we can take $\rho$ arbitrarily. By (10), we can see that $\rho$ satisfies $\mathfrak{G}$-law invariance, and it further satisfies property (a) if $\theta_{0}=\theta_{1}=\dots=\theta_{n}$. An explicit convex risk measure on $L^{\infty}$ satisfying (10) is presented in the next example, where the sources are independent.
In Example 7, even if model uncertainty exists in each source $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$, the risk measure is law invariant on each subspace as a result of robustification, discussed by Hansen and Sargent (2001).
Example 8.
The risk measure $\rho$ in (10) on $L^{\infty}(\mathcal{G}_{i})$ can be written as an entropic risk measure (Föllmer and Schied (2002)) applied to $\ell(X)$. The entropic risk measure $\mathrm{ER}_{\beta}:L^{\infty}\to\mathbb{R}$ with parameter $\beta\geq 0$ is defined by
$\mathrm{ER}_{\beta}(X)=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}\right],\text{\leavevmode\nobreak\ for\leavevmode\nobreak\ $\beta>0$},\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ $\mathrm{ER}_{0}=\mathbb{E}$}.$
The convex risk measures ER have unique properties among monetary risk measures, characterized by Embrechts et al. (2021), and they play a fundamental role in the recent characterization of monotone additive statistics by Mu et al. (2024). Note that (10) reads as $\rho(X)=\mathrm{ER}_{\theta_{i}}(\ell(X))$ on $L^{\infty}(\mathcal{G}_{i})$ for each $i\in\{0,1,\dots,n\}$. Suppose that $\mathfrak{G}=\{\mathcal{G}_{0},\mathcal{G}_{1},\dots,\mathcal{G}_{n}\}\subseteq\Sigma(\mathcal{F})$ is a collection of pair-wise independent $\sigma$-algebras. Define $\widehat{\rho}:L^{\infty}\to\mathbb{R}$ by
$\widehat{\rho}(X)=\max_{i\in\{0,1,\dots,n\}}\mathrm{ER}_{\theta_{i}}(\mathbb{E}[\ell(X)|\mathcal{G}_{i}]).$
By Jensen’s inequality, $\mathrm{ER}_{\theta_{i}}(\ell(X))\geq\mathrm{ER}_{\theta_{i}}(\mathbb{E}[\ell(X)])=\mathbb{E}[\ell(X)]$ for all $X\in L^{\infty}$ and $i\in\{0,1,\dots,n\}$. Hence, $\widehat{\rho}(X)=\mathrm{ER}_{\theta_{i}}(\ell(X))$ for all $X\in L^{\infty}(\mathcal{G}_{i})$ and thus it satisfies (10). If $\ell(x)=x$, then $\widehat{\rho}$ is a convex risk measure because it is the maximum of several convex risk measures of the form $\mathrm{ER}_{\theta_{i}}(\mathbb{E}[X|\mathcal{G}_{i}])$.
5 Extensions
In this section, we present three extensions of partially law-invaraint coherent risk measures. First, in Section 5.1, we introduce a technical tool called coherent adjustments, which admit interesting mathematical properties. Coherent adjustments allow us to reformulate Theorem 1 and serve as the foundation for the subsequent two extensions. Second, in Section 5.2, we strengthen partial law invariance to a property, which still generalizes law invariance, that allows us to extend the representation of Kusuoka (2001). Finally, in Section 5.3, we extend some of the results regarding partially law-invariant coherent risk measures to partially law-invariant convex risk measures.
5.1 Coherent adjustments and an alternative representation
Next, we take a different route to introduce the concept of coherent adjustments, a useful technical tool that helps us further understand $\mathcal{G}$-law-invariant coherent risk measures. In particular, coherent adjustments make precise how the risk measures with supporting sets $\mathcal{S}(\mu)$ are built in (5).
Definition 1.
Let $\mu\in\mathcal{M}_{1}(\mathcal{G})$, a $\mu$-coherent adjustment is a mapping $\tau:L^{\infty}\to\mathbb{R}$ satisfying
  • $\mathcal{G}$-invariance: $\tau(X+Y)=\tau(X)$ for all $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$;
  • $\mu$-monotonicity: $\mathbb{E}^{\mu}[X]+\tau(X)\leq\mathbb{E}^{\mu}[Y]+\tau(Y)$ for all $X,Y\in L^{\infty}$ with $X\leq Y$;
  • Convexity: $\tau(\lambda X+(1-\lambda)Y)\leq\lambda\tau(X)+(1-\lambda)\tau(Y)$ for all $X,Y\in L^{\infty}$ and $\lambda\in[0,1]$;
  • Positive homogeneity: $\tau(\lambda X)=\lambda\tau(X)$ for all $X\in L^{\infty}$ and $\lambda\geq 0$.
Moreover, a $\mu$-convex adjustment is a mapping that satisfies the first three properties in the above list, or $\tau=-\infty$ on $L^{\infty}$. Given $\mu\in\mathcal{M}_{1}(\mathcal{G})$ we denote the set of all Fatou $\mu$-coherent adjustments by $\mathrm{CA}(\mu)$. Let $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$, we say that $(\tau_{\mu})_{\mu\in\mathcal{R}}$ is a CA-assignment if $\tau_{\mu}\in\mathrm{CA}(\mu)$ for all $\mu\in\mathcal{R}$.
Let $\mu\in\mathcal{M}_{1}(\mathcal{G})$, from the above properties, it is straightforward that the supremum of $\mu$-coherent adjustments is again a $\mu$-coherent adjustment. A simple example of a Fatou $\mu$-coherent adjustment is $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ for any $\nu\in\mathcal{E}(\mu)$, which is easy to check. In Section E.1, Proposition 17, shows that, indeed, the supremum of the form $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ contains all possible Fatou $\mu$-coherent adjustments.
We also note that $\mathcal{G}$-invariance of $\tau$ is equivalent to $\tau(X)=0$ for all $X\in L^{\infty}(\mathcal{G})$ in the presence of subadditivity.
Remark 2.
The properties of $\mu$-coherent adjustment $\tau$ appear similar to a generalized deviation measure $D$ of Rockafellar et al. (2006), which is defined by four properties: location invariance (i.e., $D(X+c)=D(X)$ for $c\in\mathbb{R}$), subadditivity, positive homogeneity and $D(X)\geq 0$ for all $X$ with strict $D(X)>0$ when $X$ is not constant. Deviation measures naturally appear in a decomposition of insurance premiums; see Nendel et al. (2021). Certainly, location invariance is not satisfied by any monetary risk measure because it conflicts cash invariance. By choosing $\mathcal{G}=\{\varnothing,\Omega\}$ (in which case $\mathcal{M}_{1}(\mathcal{G})=\{\mu\}$ and $D_{\mu}=1=\mathbb{E}[D_{\nu}|\mathcal{G}]$ for all $\nu\in\mathcal{M}_{1}$), $\mathcal{G}$-invariance is the same as location invariance; in this case, $\tau$ and $D$ share three properties. However, there is an important difference between $\tau$ and $D$, as $\tau$ may take negative values. For instance, $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ defines a $\mu$-coherent adjustment, but not a generalized deviation measure.
The name “$\mu$-coherent adjustment” is explained by Proposition 8.
Proposition 8.
Let $\rho:L^{\infty}\to\mathbb{R}$ and $\mu\in\mathcal{M}_{1}(\mathcal{G})$. The following are equivalent.
  1. (i)
    The mapping $\rho$ is a Fatou coherent risk measure with $\rho|_{L^{\infty}(\mathcal{G})}=\mathbb{E}^{\mu}|_{L^{\infty}(\mathcal{G})}$;
  2. (ii)
    there exists $\tau\in\mathrm{CA}(\mu)$ such that $\rho=\mathbb{E}^{\mu}+\tau$.
If we assume additivity on $L^{\infty}(\mathcal{G})$ and $\mathcal{G}$-law invariance for a Fatou coherent risk measure $\rho$, then $\rho$ has a form $\mathbb{E}^{\mathbb{P}}+\tau$ for $\tau\in\mathrm{CA}(\mathbb{P}|_{\mathcal{G}})$ that is similar to Proposition 8 (ii); see Section E.2.
For some $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, by Proposition 8, the following risk measure
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$ (11)
is coherent, as it is the supremum of coherent risk measures. Theorem 3 below shows that this mapping is Fatou continuous and that all Fatou coherent risk measures have this representation.
To explain why $\mu$-coherent adjustments help us extend a coherent risk measure on $L^{\infty}(\mathcal{G})$ to $L^{\infty}$, let $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ be a law-invariant coherent risk measure. For a convex $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-dense in $\mathcal{S}_{\tilde{\rho}}$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, let $\rho$ be given by (11). It is easy to see that $\rho|_{L^{\infty}(\mathcal{G})}=\tilde{\rho}$, meaning that it is $\mathcal{G}$-law invariant. In the special case that $\tau_{\mu}=0$ for all $\mu\in\mathcal{R}$, this extension technique precisely gives the class studied in Proposition 6, because $\widehat{\mathcal{S}}_{\rho}=\widehat{\mathcal{S}}_{\tilde{\rho}}\subseteq L^{1}(\mathcal{G})$.
The above ideas lead to the following alternative representation of partially law-invariant and Fatou coherent risk measures.
Theorem 3.
The mapping $\rho:L^{\infty}\to\mathbb{R}$ is a Fatou coherent risk measure if and only if there exists a convex set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$ (12)
Moreover, in (12), $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{R}$ is almost $\mathcal{G}$-law invariant.
Remark 3.
There is a straightforward way to switch between representations (4) and (12) for a Fatou and $\mathcal{G}$-law-invariant coherent risk measure. Given a convex set $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that represents $\rho$ in Theorem 1, define $\mathcal{R}=\mathcal{S}^{\mathcal{G}}$, which is almost $\mathcal{G}$-law invariant. For each $\mu\in\mathcal{R}$, define $\mathcal{S}(\mu)=\mathcal{E}(\mu)\cap\mathcal{S}$ and $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\}$. These can be used in representation (12). Conversely, let an almost $\mathcal{G}$-law invariant convex set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ representing $\rho$ in Theorem 3 be given. For $\mu\in\mathcal{R}$, let $\mathcal{S}_{\tau_{\mu}}$ be the $w(\mathcal{M},L^{\infty})$-closed and convex set in Proposition 17 (see Section E.1) that represents $\tau_{\mu}$. It is easy to show that $\mathcal{S}=\bigcup_{\mu\in\mathcal{R}}\mathcal{S}_{\tau_{\mu}}$ is convex and $S^{\mathcal{G}}=\mathcal{R}$. We can use $\mathcal{S}$ in representation (4). The above connection clarifies the construction in (5).
Uniqueness of representation (12) under a stronger notion of continuity is addressed in Section E.3. Representation (12) has little resemblance to the classical representation result of Kusuoka (2001) for law-invariant coherent risk measures. The following section addresses this gap by further imposing a necessary and sufficient condition for a similar representation along that direction.
5.2 Strong partial law invariance and Kusuoka-type representation
Next, we introduce a new property, which is stronger than partial law invariance and is necessary and sufficient to reduce (12) to a Kusuoka-type representation; see Proposition 11 in Appendix A.
Define $\ker(\mathcal{G})=\{X\in L^{\infty}:\mathbb{E}[X|\mathcal{G}]=0\}.$ We say that $\rho:L^{\infty}\to\mathbb{R}$ is strongly $\mathcal{G}$-law invariant if for all $Z\in\ker(\mathcal{G})$ and $X,Y\in L^{\infty}(\mathcal{G})$ with $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$ we have
$\rho(Z+X)=\rho(Z+Y).$
Taking $Z=0\in\ker(\mathcal{G})$, we see that strong $\mathcal{G}$-law invariance implies $\mathcal{G}$-law invariance, hence the name. Since $\ker(\mathcal{F})=\{0\}$, the risk measure $\rho$ is law invariant if and only if $\rho$ is strongly $\mathcal{F}$-law invariant. Therefore, strong $\mathcal{G}$-law invariance and $\mathcal{G}$-law invariance are both a generalization of law invariance. The following result gives an equivalent condition to strong $\mathcal{G}$-law invariance.
Lemma 1.
The mapping $\rho$ is strongly $\mathcal{G}$-law invariant if and only if for all $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$ with $Y\overset{\mathrm{d}}{=}_{\mathbb{P}}\mathbb{E}[X|\mathcal{G}]$ we have
$\rho(X)=\rho(X-\mathbb{E}[X|\mathcal{G}]+Y).$
An example of strongly $\mathcal{G}$-law invariant risk measure $\rho$ is $\rho(X)=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$ given in (7), where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is law invariant. Examples of functionals satisfying $\mathcal{G}$-law invariance but not strong $\mathcal{G}$-law invariance will be provided in Section E.5.
Before we give the main result of the section, we need some technical details. Given $X\in L^{1}$, $\mu\in\mathcal{M}_{1}$ and $\alpha\in[0,1)$, define $F_{\mu,X}^{-1}(\alpha)=\inf\{x\in\mathbb{R}\mid F_{\mu,X}(x)\geq\alpha\}$ and $q_{X}(\alpha)=F_{\mathbb{P},X}^{-1}(\alpha)$. Given $\mu\in\mathcal{M}_{1}(\mathcal{G})$, we can define $Q\in\mathcal{M}_{B}$ (see Föllmer and Schied (2016, Theorem 4.62)) by
$Q([0,t))=\int_{[0,t)}(1-s)\leavevmode\nobreak\ \mathrm{d}q_{D_{\mu}}(s),\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }t\in[0,1).$ (13)
[…]
$\sup_{\mu\in\mathcal{M}_{Q}}\mathbb{E}^{\mu}[X]=\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])\leavevmode\nobreak\ Q(\mathrm{d}\alpha);$ (14)
see Föllmer and Schied (2016, Theorem 4.62). Let $\mathrm{CA}(Q)=\bigcap_{\mu\in\mathcal{M}_{Q}}\mathrm{CA}(\mu).$ This set is non-empty since it contains the zero map. We present the following result, generalizing the representation of Kusuoka (2001).
Theorem 4.
Let $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
    The mapping $\rho$ is a Fatou and strongly $\mathcal{G}$-law-invariant coherent risk measure;
  2. (ii)
    there exist $\mathcal{D}\subseteq\mathcal{M}_{B}$ and $\tau_{Q}\in\mathrm{CA}(Q)$ for each $Q\in\mathcal{D}$ such that
    $\rho(X)=\sup_{Q\in\mathcal{D}}\left\{\int_{[0,1)}\mathrm{ES}_{\alpha}\left(\mathbb{E}[X|\mathcal{G}]\right)\leavevmode\nobreak\ Q(\mathrm{d}\alpha)+\tau_{Q}(X)\right\}.$ (15)
In the setting $\mathcal{G}=\mathcal{F}$, the representation (15) reduces to the Kusuoka representation,
$\rho(X)=\sup_{Q\in\mathcal{D}}\int_{[0,1)}\mathrm{ES}_{\alpha}(X)\leavevmode\nobreak\ Q(\mathrm{d}\alpha),\hskip 3.0ptX\in L^{\infty},$
because in this case $\mathbb{E}[X|\mathcal{G}]=X$, $\tau_{Q}(X)=0$, and strong $\mathcal{G}$-law invariance is law invariance. A simple example is to choose $\tau_{Q}=0$ for all $Q\in\mathcal{D}$, which yields the risk measure $\rho:X\mapsto\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$ studied in Proposition 6.
5.3 Convex risk measures
This section discusses convex risk measures and their relation to the variational preferences model of Maccheroni et al. (2006). We provide representations for partially law-invariant convex risk measures akin to Theorem 3. A risk measure $\rho:L^{\infty}\to\mathbb{R}$ is Fatou and convex if and only if there exists $\alpha:\mathcal{M}_{1}\to(-\infty,\infty]$ such that
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}(\mathbb{E}^{\mu}[X]-\alpha(\mu)),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$
see Föllmer and Schied (2016). Using the subjective loss interpretation for the random variables, we see that applying convex risk measures to subjective losses agrees with the variational preferences model of Maccheroni et al. (2006) in the case that $\inf_{\mu\in\mathcal{M}_{1}}\alpha(\mu)=0$.
Similar to CA-assignments, we say that $(\tau_{\mu})_{\mu\in\mathcal{M}_{1}(\mathcal{G})}$ is a CoA-assignment if for all $\mu\in\mathcal{M}_{1}(\mathcal{G})$, $\tau_{\mu}$ is a Fatou $\mu$-convex adjustment and the mapping $\mu\mapsto\tau_{\mu}(0)$ is concave. The following result is the convex version of Theorem 3.
Theorem 5.
The mapping $\rho:L^{\infty}\to\mathbb{R}$ is a Fatou convex risk measure if and only if there exists a CoA-assignment $(\tau_{\mu})_{\mu\in\mathcal{M}_{1}(\mathcal{G})}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Moreover, $\rho$ is $\mathcal{G}$-law invariant if and only if for all $c>-\rho(0)$ the set $\left\{\mu\in\mathcal{M}_{1}(\mathcal{G})\mid-\tau_{\mu}(0)\leq c\right\}$ is almost $\mathcal{G}$-law invariant.
The risk measures ER in Example 8 are the most common examples of convex risk measures. We next discuss a few $\mathcal{G}$-law-invariant versions of ER, which shed some light on different ways to construct partially law-invariant convex risk measures. In the examples below, partial law invariance is not necessarily associated with the interpretation of model uncertainty.
We write the conditional version of ER (this is not a risk measure in our definition) as
$\mathrm{ER}_{\beta}(X|\mathcal{G})=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
There are many ways to generalize the family of ER to satisfy $\mathcal{G}$-partial law-invariance, and we list three simple choices. The first is to use the mean of the conditional ER, that is,
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathbb{E}[\mathrm{ER}_{\beta}(X|\mathcal{G})]=\mathbb{E}\left[\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
The second is to compute the ER of the conditional mean in the form of (7), that is,
$\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathrm{ER}_{\beta}(\mathbb{E}[X|\mathcal{G}])=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta\mathbb{E}[X|\mathcal{G}]}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Finally, we can also compute ER of the conditional ER, but this gives back the usual ER, that is
$\mathrm{ER}_{\beta}(\mathrm{ER}_{\beta}(X|\mathcal{G}))=\frac{1}{\beta}\log\mathbb{E}\left[\exp\left(\beta\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right)\right]=\mathrm{ER}_{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
(This is time-consistency of ER as a dynamic risk measure; see Kupper and Schachermayer (2009).) The last version is fully law invariant and not interesting to us. The first two versions are $\mathcal{G}$-law invariant and have intriguing and distinct properties. For instance, if $X$ is $\mathcal{G}$-measurable, then
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathbb{E}[X]\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathrm{ER}_{\beta}(X);$
if $Y$ is independent of $\mathcal{G}$, then
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)=\mathrm{ER}_{\beta}(Y)\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)=\mathbb{E}[Y].$
The above formulas show that ${\mathrm{ER}}^{\mathcal{G}}_{\beta}$ and $\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}$ are not fully law invariant, and none of them dominates the other. Moreover, if $X$ and $Y$ above are identically distributed and non-degenerate, then ${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)<{\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)$ and $\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)>\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)$. In other words, ${\mathrm{ER}}^{\mathcal{G}}_{\beta}$ penalizes risks independent of $\mathcal{G}$, whereas $\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}$ penalizes risks measurable to $\mathcal{G}$. This shows different emphases of the two risk measures when assessing risks from different sources; note that here we do not interpret events outside $\mathcal{G}$ as events with ambiguity; indeed, all computations are carried with $\mathbb{P}$ that is known and fixed in this setting.
6 Partially law-invariant ES and numerical illustrations
This section gives tractable formulas for an important class of partially law-invariant risk measures, the $\mathcal{G}$-law invariant ES introduced in Section 2.2, given by
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$ and satisfies some regularity (to be made rigorous later). Afterward, we numerically implement the section’s results and see how decisions in a simple investment vary as we change the parameters in the risk measure.
We focus on ES here for two reasons. First, ES is the standard regulatory risk measure in the banking sector (BCBS (2019)), and hence the most relevant for financial applications. Second, ES has special advantages in optimization (Rockafellar and Uryasev (2002)), and the partially law-invariant version of ES admits tractable formulas (Proposition 9 below).
To prepare for the formula of $\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}$, we first define two technical terms.
Definition 2.
Let $\mathcal{X}\subseteq L^{\infty}$, we call $Y\in L^{\infty}$ the essential supremum of $\mathcal{X}$ if
  1. (i)
    For all $X\in\mathcal{X}$, $X\leq Y$;
  2. (ii)
    if $Z\in L^{\infty}$ and $X\leq Z$ for all $X\in\mathcal{X}$, then $Y\leq Z$.
We will denote $Y$ by $\operatorname*{ess\text{-}sup}\mathcal{X}$, as the essential supremum (if it exists) is $\mathbb{P}$-a.s. unique.
Definition 3.
We say that $\mathcal{X}\subseteq L^{1}$ is $\mathcal{G}$-convex if for all $X,Y\in\mathcal{X}$ and $\lambda\in L^{\infty}(\mathcal{G})$ with $0\leq\lambda\leq 1$, we have $\lambda X+(1-\lambda)Y\in\mathcal{X}.$ We say that $\mathcal{R}\subseteq\mathcal{M}$ is $\mathcal{G}$-convex if $\widehat{\mathcal{R}}$ is $\mathcal{G}$-convex.
Given certain conditions on $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$, the $\mathcal{G}$-law invariant risk measure $\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}$ mentioned in Section 2.2 can be determined via a minimization formula, similar to the formulation of ES by Rockafellar and Uryasev (2002). Below, a coherent conditional risk measure is a mapping from $L^{\infty}$ to $L^{\infty}(\mathcal{G})$ satisfying the for axioms of coherence but with constants replaced by elements of $L^{\infty}(\mathcal{G})$; see Detlefsen and Scandolo (2005).
Proposition 9.
Let $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$ be $w(\mathcal{M},L^{\infty})$-compact and $\mathcal{G}$-convex. For $\alpha\in[0,1)$, we have
$\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}\left[\rho_{\mathcal{G}}^{\mathcal{Q}}\left((X-x)^{+}\right)\right]\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\rho_{\mathcal{G}}^{\mathcal{Q}}:L^{\infty}\to L^{\infty}(\mathcal{G})$ is the coherent conditional risk measure
$\rho_{\mathcal{G}}^{\mathcal{Q}}(X)=\operatorname*{ess\text{-}sup}_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X|\mathcal{G}],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Moreover, the set $\mathcal{Q}_{\beta}=\left\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\right\}$ for $\beta\in[0,1)$ is $w(\mathcal{M},L^{\infty})$-compact and $\mathcal{G}$-convex.
Denote by $\mathrm{ES}_{\beta}(X|\mathcal{G})=\rho_{\mathcal{G}}^{\mathcal{Q}_{\beta}}(X)$, $X\in L^{\infty}$, which is the conditional version of ES in Detlefsen and Scandolo (2005). The benefit of using $\mathcal{Q}_{\beta}$ is that under tractable assumptions of the conditional distribution of $X$ given $\mathcal{G}$, $\mathbb{E}[\mathrm{ES}_{\beta}((X-x)^{+}|\mathcal{G})]$ can be computed using conditional distributions.
Next, we illustrate with a Gaussian example in the context of investment. Let $(\Omega,\mathcal{F})=(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$ and $\mathbb{P}$ be multivariate Gaussian with mean $m\in\mathbb{R}^{2}$ and covariance matrix $\Sigma\in\mathbb{R}^{2\times 2}$, where $\Sigma_{ij}=(c+(1-c)\mathds{1}_{\{i=j\}})\sigma_{i}\sigma_{j}$ for $i,j\in\{1,2\}$. Let $X_{1},X_{2}$ be the random variables given by the projections. We interpret $X_{1},X_{2}$ as the losses (as a percentage) of two positions on a market and $\mathbb{P}$ as their estimated bivariate distribution. Define $\mathcal{G}=\sigma(X_{1})$, the goal is to use the risk measure $\rho_{\beta}=\sup_{\mu\in\mathcal{Q}^{\beta}}\mathrm{ES}^{\mu}_{\alpha}$ to measure the risk between different portfolios given by $\pi_{1}X_{1}+\pi_{2}X_{2}$, where $(\pi_{1},\pi_{2})$ is a vector in the unit simplex in $\mathbb{R}^{2}$.
In Appendix F, for $\beta\in[0,1)$, we derive an integral formula for
$f_{\beta}(\pi_{1},\pi_{2},x)=\mathbb{E}\left[\mathrm{ES}_{\beta}\left((\pi_{1}X_{1}+\pi_{2}X_{2}-x)^{+}|X_{1}\right)\right],$ (16)
which we can compute via numerical integration. Therefore, by Proposition 9,
$\overline{\mathrm{ES}}_{\alpha}^{\mathcal{Q}_{\beta}}\left(\pi_{1}X_{1}+\pi_{2}X_{2}\right)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}f_{\beta}(\pi_{1},\pi_{2},x)\right),$
which we can compute via convex optimization.
Figures 1 and 2 display plots depicting the risk, calculated using $\rho_{\beta}$, associated with various portfolios. The horizontal axes in all plots represent the weight allocated to $X_{1}$. In all cases, $\rho_{\beta}(X_{1})=\mathrm{ES}_{\alpha}(X_{1})$, explaining why all the lines agree when $\pi_{1}=1$.
Figure 1 aligns with our theoretical expectations. Specifically, when $m_{2}=0$ (plot 1 in Figure 1), an agent who assesses their risk using $\beta=0$ opts for a balanced investment in $X_{1}$ and $X_{2}$. This conclusion arises since the risk measure $\rho_{0}$ coincides with $\mathrm{ES}_{\alpha}$, and the agent attains the lowest risk through total diversification. As $\beta$ increases, the optimally risk-minimized portfolio progressively shifts towards an investment completely in $X_{1}$. This happens because the agent is becoming increasingly fearful that the distribution of $X_{2}$ is misspecified. The additional plots within Figure 1 explore the scenarios when $m_{2}=-0.05,-0.1$, indicating instances when $X_{2}$ is presumed to yield a positive average return. The observed trend mirrors the behavior noted in the scenario where $m_{2}=0$, albeit with reduced intensity.
Figure 1: Varying $\beta$, $m_{1}=0$, $\sigma_{1}=0.1$, $\sigma_{2}=0.1$, $c=0.5$, $\alpha=0.95$; the red dotted curve connects the optimizers for different $\beta$
Figure 2 fixes $\beta=0.95$ but varies $\sigma_{2}$. In all plots, an agent who assesses their risk using $\rho_{0.95}$ invests fully into $X_{2}$ when the standard deviation of $X_{2}$ is low enough. The agent fears model misspecification, with the parameter $\beta$ quantifying this fear. Since $\beta=0.95$, the agent heavily suspects that $X_{2}$ is misspecified. However, when $\sigma_{2}$ is very small, because the distributions in “the $\beta$-neighbourhood” of the estimated distribution of $X_{2}$ are considerably more favorable than the estimated distribution of $X_{1}$, the agent still opts for a complete investment in $X_{2}$ despite the fear of misspecification.
Figure 2: Varying $\sigma_{2}$, $m_{1}=0$, $\sigma_{1}=0.1$, $c=0.5$, $\alpha=0.95$, $\beta=0.95$; the red dotted curve connects the optimizers for different $\sigma_{2}$
7 Conclusion
Our main purpose is to propose and study the concept of partial law invariance as a generalization of law invariance for risk measures and preference relations. This new property is appealing from a decision-theoretic perspective and interesting from a purely mathematical perspective. To summarize the main motivation, different economic sources have different levels of model uncertainty, and a law-invariant mapping or preference relation does not take these differences into account, thus calling for the concept of partial law invariance to capture this discrepancy.
The economic relevance of the new framework is illustrated throughout the paper with examples. In addition to conceptual development, the new framework also leads to mathematical novelty in terms of new classes of risk measures and new representation results in forms not present in the literature. In particular, coherent and convex adjustments play an important role in understanding the new property and its representation in both the context of coherent and convex risk measures.
The possibility of characterizing partial law invariance to other classes of risk measures may be worth discussing. The star-shaped risk measures, studied by Castagnoli et al. (2022), are a natural generalization of convex risk measures, and they satisfy the following property in addition to $\rho(0)=0$.
  1. Star-shapedness: $\rho(\lambda X)\leq\lambda\rho(X)$ for all $X\in L^{\infty}$ and $\lambda\in[0,1]$.
Representations of law-invariant (quasi-)star-shaped risk measures have recently been studied by Han et al. (2021) and Laeven et al. (2024), which admit a quite different structure than the Kusuoka-type. Another research question is how to incorporate different probabilities, instead of projections of the same probability $\mathbb{P}$ as in our setting, in the context of multiple sources.
Acknowledgements
The authors thank Peter Wakker for introducing the source theory to RW, which inspired the idea of this paper, and for many helpful comments on a previous version of the paper. The authors also thank Felix Liebrich and Frank Riedel for their excellent comments that led to substantial improvements in the paper. YS and RW are financially supported by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2020-04356, RGPIN-2024-03728, CRC-2022-00141).
Appendix A Review on risk measures
This appendix reviews risk measures in the framework of Artzner et al. (1999).
We say that a risk measure $\rho$ is Lebesgue (continuous) if for all uniformly bounded sequences $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ that converge $\mathbb{P}$-a.s. to some $X\in L^{\infty}$, it holds that
$\lim_{n\to\infty}\rho(X_{n})=\rho(X).$
An example of a Lebesgue (thus Fatou) coherent risk measure is $\mathrm{ES}_{\alpha}^{\mu}$ for $\mu\in\mathcal{M}_{1}$ and $\alpha\in[0,1)$.
Some well-known facts are collected below. A coherent risk measure $\rho$ is Fatou if and only if $\rho$ is $w(L^{\infty},\mathcal{M})$-lower semicontinuous, a proof of which can be found in Föllmer and Schied (2016, Theorem 4.33). A coherent risk measure has a dual representation in the following form; see, e.g., Föllmer and Schied (2016, Theorem 4.33). Recall that for a Fatou coherent risk measure $\rho:L^{\infty}\to\mathbb{R}$, it’s supporting set is given by
$\mathcal{S}_{\rho}=\left\{\mu\in\mathcal{M}_{1}:\mathbb{E}^{\mu}[X]\leq\rho(X)\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ X\in L^{\infty}\right\}.$
Proposition 10 (Delbaen (2002)).
For $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
    The mapping $\rho$ is a Fatou coherent risk measure;
  2. (ii)
    there exists a $w(\mathcal{M},L^{\infty})$-closed and convex set $\mathcal{S}\subseteq\mathcal{M}_{1}$ such that
    $\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$ (17)
In case (ii) holds, $\mathcal{S}$ is uniquely given by $\mathcal{S}_{\rho}$ (the supporting set of $\rho$). Moreover, $\mathcal{R}\subseteq\mathcal{M}_{1}$ gives representation (17) if and only if $\operatorname*{conv}(\mathcal{R})$ is $w(\mathcal{M},L^{\infty})$-dense in $\mathcal{S}_{\rho}$.
Example 9.
For $\mu\in\mathcal{M}_{1}$ and $\alpha\in[0,1)$, $\mathrm{ES}^{\mu}_{\alpha}$ admits the dual representation
$\mathrm{ES}^{\mu}_{\alpha}(X)=\sup_{\nu\in\mathcal{S}_{\alpha,\mu}}\mathbb{E}^{\nu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\mathcal{S}_{\alpha,\mu}=\{\nu\in\mathcal{M}_{1}\mid({1-\alpha})D_{\nu}\leq D_{\mu}\}$; see e.g., Föllmer and Schied (2016, Theorem 4.52). Since $\mathcal{S}_{\alpha,\mu}$ is $w\left(\mathcal{M},L^{\infty}\right)$-closed and convex, $\mathcal{S}_{\mathrm{ES}^{\mu}_{\alpha}}=\mathcal{S}_{\alpha,\mu}$ holds.
Law-invariant coherent risk measures also admit a representation of Kusuoka (2001), and Fatou continuity is automatic for law-invariant convex risk measures, as shown by Jouini et al. (2006) and Svindland (2010). Let $\mathcal{M}_{B}$ denote the probability measures on $[0,1)$ equipped with the restricted Borel $\sigma$-algebra.
Proposition 11 (Kusuoka (2001); Jouini et al. (2006); Svindland (2010)).
For $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
    The mapping $\rho$ is a Fatou law-invariant coherent risk measure;
  2. (ii)
    the mapping $\rho$ is a law-invariant coherent risk measure;
  3. (iii)
    there exists some $\mathcal{D}\subseteq\mathcal{M}_{B}$ such that
    $\rho(X)=\sup_{Q\in\mathcal{D}}\int_{[0,1)}\mathrm{ES}_{\alpha}(X)\leavevmode\nobreak\ Q(\mathrm{d}\alpha),\hskip 3.0ptX\in L^{\infty}.$
It is well known that the monetary risk measure $\rho:L^{\infty}\to\mathbb{R}$ is convex if and only if $\mathcal{A}_{\rho}$ is convex, and coherent if and only if $\mathcal{A}_{\rho}$ is a convex cone.
Appendix B Additional results and proofs accompanying Section 2
Proof of Proposition 1.
[…]
Proof of Proposition 2.
[…]
Proof of Proposition 3.
Let $X,Y\in L^{\infty}(\mathcal{G})$ and $X\buildrel\mathrm{d}\over{=}_{\mathbb{P}}Y$. For $\mu\in\mathcal{Q}$ and $x\in\mathbb{R}$ we have
$\mu(X\leq x)=\mathbb{E}\left[D_{\mu}\mathds{1}_{\{X\leq x\}}\right]=\mathbb{E}\left[D_{\mu}\right]\mathbb{E}\left[\mathds{1}_{\{X\leq x\}}\right]=\mathbb{P}(X\leq x)$
and similarly $\mu(Y\leq x)=\mathbb{P}(Y\leq x)$. Hence, $X$ and $Y$ are identically distributed under each element of $\mathcal{Q}$. Since $\rho$ is $\mathcal{Q}$-based, we have $\rho(X)=\rho(Y)$. This implies that $\rho$ is $\mathcal{G}$-law invariant. ∎
The next remark discusses how Proposition 3 is related to conditional non-atomicity proposed by Shen et al. (2019).
Remark 4.
Let $\mathcal{Q}$ be a collection of probability measures on $(\Omega,\mathcal{F})$ and $\mathcal{M}_{1}(\mathcal{Q})$ be the set of probability measures dominating (in the sense of absolute continuity) each element of $\mathcal{Q}$, and we assume it is non-empty (it is always non-empty if $\mathcal{Q}$ is finite or $\mathcal{Q}\subseteq\mathcal{M}_{1}$). The set $\mathcal{Q}$ is said to be conditionally atomless if there exist $\nu\in\mathcal{M}_{1}(\mathcal{Q})$ and random variable $X$ such that under $\nu$, $X$ is continuously distributed and independent of $\sigma(D_{\mu}:\mu\in\mathcal{Q})$. As discussed by Delbaen (2021), an equivalent statement for conditional non-atomicity of $\mathcal{Q}$ is that there exists an atomless $\sigma$-algebra $\mathcal{G}$ independent of $\sigma(D_{\mu}:\mu\in\mathcal{Q})$ under some $\nu\in\mathcal{M}_{1}(\mathcal{Q})$. Clearly, when $(\Omega,\mathcal{G},\mathbb{P})$ is atomless, the condition that $\mathcal{G}$ is independent of $\sigma(D_{\mu}:\mu\in\mathcal{Q})$ in Proposition 3 is sufficient for conditional non-atomicity of $\mathcal{Q}$.
Appendix C Additional results and proofs accompanying Section 3
C.1 Additional results and proofs accompanying Section 3.1
[…]
  1. Translation invariance: $X+c\succsim Y+c.$ for all $X,Y\in L^{\infty}$ with $X\succsim Y$ and $c\in\mathbb{R}$;
  2. Convexity: $\lambda X+(1-\lambda)Y\succsim Z$ for all $X,Y,Z\in L^{\infty}$ with $X\succsim Z$ and $Y\succsim Z$ and $\lambda\in[0,1]$;
  3. Positive homogeneity: $\lambda X\succsim\lambda Y$ for all $X,Y\in L^{\infty}$ with $X\succsim Y$ and $\lambda\geq 0$;
  4. Continuity from below: for all increasing sequences $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ that converges to $X$ and $Y\in L^{\infty}$ with $X_{n}\succsim Y$ for $n\in\mathbb{N}$, we have $X\succsim Y$.
The next proposition clarifies the connection between CRM preferences and these properties, and we provide a simple self-contained proof.
Proposition 12.
[…]
  1. (i)
[…]
  2. (ii)
[…]
Proof.
[…]
$\rho(X)\leq l=\lim_{k\to\infty}\rho(X_{k})\leq\rho(X),$
which implies $\rho(X_{n})\uparrow\rho(X)$. For the converse, assume that $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ is an increasing sequence that converges to $X$ and $Y\in L^{\infty}$ with $X_{n}\succsim Y$ for $n\in\mathbb{N}$. Therefore, it holds that $\rho(X_{n})\leq\rho(Y)$ for all $n\in N$ and
$\rho(X)=\lim_{n\to\infty}\rho(X_{n})\leq\rho(Y),$
meaning that $X\succsim Y$.
As $\rho$ is monetary, it is a well-known fact that $\rho$ is Fatou continuous if and only if for all increasing sequences $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ that converges to $X$, we have $\rho(X_{n})\uparrow\rho(X)$, which proves the claim. ∎
Proof of Proposition 4.
[…]
Proof of Proposition 5.
The forward is trivial. For the converse note that $\{X\in L^{\infty}\mid\rho(X)\leq c\}=\mathcal{A}_{\rho}+c$. Therefore, the level sets of $\rho$ are $\mathcal{G}$-law invariant; the claim is straightforward to prove after this observation. ∎
C.2 Discussion and proof of Theorem 1
A proof of Theorem 1 follows from an analysis of the supporting set of $\rho$, which will be presented in a few separate results.
Proposition 13.
Let $\rho:L^{\infty}\to\mathbb{R}$ be a Fatou coherent risk measure and $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$. Then, $\tilde{\rho}$ is a Fatou coherent risk measure. Furthermore, $\mathcal{S}_{\rho}^{\mathcal{G}}\subseteq\mathcal{S}_{\tilde{\rho}}$ and $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-dense in $\mathcal{S}_{\tilde{\rho}}$.
Proof.
The fact that $\tilde{\rho}$ is a Fatou coherent risk measure is clear. Since $\mathcal{S}_{\rho}^{\mathcal{G}}$ is convex and for $X\in L^{\infty}(\mathcal{G})$,
$\tilde{\rho}(X)=\rho(X)=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}^{\mu}[X]=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[\mathbb{E}[D_{\mu}|\mathcal{G}]X]=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\mathbb{E}^{\mu}[X],$
we have that $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-dense in $\mathcal{S}_{\tilde{\rho}}$ by Proposition 10. ∎
It is natural to wonder when $\mathcal{S}_{\rho}^{\mathcal{G}}=\mathcal{S}_{\tilde{\rho}}$ in Proposition 13. Proposition 14 offers sufficient conditions for this equality.
Lemma 2.
The mapping $R:\mathcal{M}\to\mathcal{M}(\mathcal{G}):\mu\mapsto\mu|_{\mathcal{G}}$ is $w(\mathcal{M},L^{\infty})/w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-continuous.
Proof.
Let $\mu_{\lambda},\mu\in\mathcal{M}$ with $\lim_{\lambda}\mu_{\lambda}=\mu$ in $w(\mathcal{M},L^{\infty})$. Fix some $X\in L^{\infty}(\mathcal{G})$. We have
$\mathbb{E}^{R(\mu_{\lambda})}[X]=\mathbb{E}^{\mu_{\lambda}}[X]\to\mathbb{E}^{\mu}[X]=\mathbb{E}^{R(\mu)}[X].$
Therefore $R(\mu_{\lambda})\to R(\mu)$ in $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$. ∎
Proposition 14.
Let $\rho:L^{\infty}\to\mathbb{R}$ be a Fatou coherent risk measure and $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$. The following are sufficient conditions for $\mathcal{S}_{\rho}^{\mathcal{G}}=\mathcal{S}_{\tilde{\rho}}$.
  1. (i)
    The risk measure $\rho$ satisfies $\rho(\mathbb{E}[X|\mathcal{G}])\leq\rho(X)$ for all $X\in L^{\infty}$ (equivalently, $\widehat{\mathcal{S}}^{\mathcal{G}}_{\rho}\subseteq\widehat{\mathcal{S}}_{\rho}$).
  2. (ii)
    The risk measure $\rho$ is Lebesgue continuous.
Proof.
$(\mathrm{i})$ For $\mu\in\mathcal{S}_{\tilde{\rho}}$, we have
$\mathbb{E}^{\mu}[X]=\mathbb{E}[D_{\mu}\mathbb{E}[X|\mathcal{G}]]\leq\rho(\mathbb{E}[X|\mathcal{G}])\leq\rho(X),$
for all $X\in L^{\infty}$. Therefore, the extension of $\mu$ to $\mathcal{F}$, i.e., the measure $\hat{\mu}:\mathcal{F}\to[0,1]$ given by $A\mapsto\mathbb{E}^{\mu}[\mathds{1}_{A}]$, is in $\mathcal{S}_{\rho}$. Since $\hat{\mu}|_{\mathcal{G}}=\mu$, we have $\mathcal{S}_{\tilde{\rho}}\subseteq\mathcal{S}_{\rho}^{\mathcal{G}}$. $(\mathrm{ii})$ The restriction $\mu\mapsto\mu|_{\mathcal{G}}$ is $w(\mathcal{M},L^{\infty})/w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-continuous by Lemma 2 and $\mathcal{S}_{\rho}$ is $w(\mathcal{M},L^{\infty})$-compact (see Föllmer and Schied (2016, Corollary 4.38)) we have that $\mathcal{S}^{\mathcal{G}}_{\rho}$ is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-compact. Therefore, $\mathcal{S}_{\rho}^{\mathcal{G}}$ is closed since $w(L^{1},L^{\infty})$ is Hausdorff, which implies $\mathcal{S}_{\rho}^{\mathcal{G}}=\mathcal{S}_{\tilde{\rho}}$. ∎
Remark 5.
Although Proposition 14 only provides sufficient conditions for $\mathcal{S}_{\rho}^{\mathcal{G}}=\mathcal{S}_{\tilde{\rho}}$, we did not find any counter-example in which $\mathcal{S}_{\rho}^{\mathcal{G}}\neq\mathcal{S}_{\tilde{\rho}}$, which seems to be a challenging question. This relates to whether we can omit the operation of taking a closure in Theorem 1. If $\mathcal{S}_{\rho}^{\mathcal{G}}=\mathcal{S}_{\tilde{\rho}}$ always holds, then that operation can be omitted.
We are ready to prove Theorem 1.
Proof of Theorem 1.
$(\mathrm{i})\Rightarrow(\mathrm{iii})$: Since $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$ is law-invariant. We have that $\mathcal{S}_{\tilde{\rho}}$ is $\mathcal{G}$-law invariant. By Proposition 13, $\mathcal{S}^{\mathcal{G}}_{\rho}$ is almost $\mathcal{G}$-law invariant. $(\mathrm{iii})\Rightarrow(\mathrm{ii})$: By Proposition 10, representation (4) holds with $\mathcal{S}_{\rho}$. $(\mathrm{ii})\Rightarrow(\mathrm{i})$: It is clear that $\rho$ is a Fatou coherent risk measure. It is also clear that $\mathcal{S}^{\mathcal{G}}$ is convex and supports $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$. Since $\mathcal{S}^{\mathcal{\mathcal{G}}}$ is almost $\mathcal{G}$-law invariant, $\tilde{\rho}$ is law invariant. This implies that $\rho$ is $\mathcal{G}$-law invariant. ∎
C.3 Additional results and proofs accompanying Section 3.2
Proposition 15.
Let $\tilde{\rho}$ in (7) be monotonic (resp. coherent) and Fatou. Then, the resulting risk measure $\rho$ in (7) is also monotonic (resp. coherent) and Fatou.
Proof.
It is straightforward that $\rho$ is monotonic. Let $(X_{n})_{n\in\mathbb{N}}$ be an increasing sequence in $L^{\infty}$ that converges $\mathbb{P}$-a.s. to some $X\in L^{\infty}$. Since $(X_{n})_{n\in\mathbb{N}}$ is uniformly bounded, it is uniformly integrable. Thus $X_{n}$ converges to $X$ in $L^{1}$. Therefore, $(\mathbb{E}[X_{n}|\mathcal{G}])_{n\in\mathbb{N}}$ converges in $L^{1}$ to $\mathbb{E}[X|\mathcal{G}]$ by Jensen’s inequality. Thus, $(\mathbb{E}[X_{n}|\mathcal{G}])_{n\in\mathbb{N}}$ converges in probability to $\mathbb{E}[X|\mathcal{G}]$. Since $(\mathbb{E}[X_{n}|\mathcal{G}])_{n\in\mathbb{N}}$ is also increasing, it converges $\mathbb{P}$-a.s. to $\mathbb{E}[X|\mathcal{G}]$. Therefore, by Fatou continuity of $\tilde{\rho}$,
$\lim_{n\to\infty}\rho(X_{n})\leq\rho(X)=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])\leq\liminf_{n\to\infty}\tilde{\rho}(\mathbb{E}[X_{n}|\mathcal{G}])=\lim_{n\to\infty}\rho(X_{n}).$
So $\rho(X)=\lim_{n\to\infty}\rho(X_{n})$. This means that $\rho$ is continuous from below. It is well-known that continuity from below implies Fatou-continuity under the monotonicity assumption. The fact that $\rho$ is coherent is straightforward. ∎
Proposition 16.
If $X_{\lambda}\to X$ in $w\left(L^{1},L^{\infty}\right)$ then $\mathbb{E}[X_{\lambda}|\mathcal{G}]\to\mathbb{E}[X|\mathcal{G}]$ in $w\left(L^{1},L^{\infty}\right)$.
Proof.
Let $X_{\lambda}\to X$ in $w\left(L^{1},L^{\infty}\right)$ and $Y\in L^{\infty}$, we have
$\mathbb{E}[\mathbb{E}[X_{\lambda}|\mathcal{G}]Y]=\mathbb{E}[X_{\lambda}\mathbb{E}[Y|\mathcal{G}]]\to\mathbb{E}[X\mathbb{E}[Y|\mathcal{G}]]=\mathbb{E}[\mathbb{E}[X|\mathcal{G}]Y].$
Implying that $\mathbb{E}[X_{\lambda}|\mathcal{G}]\to\mathbb{E}[X|\mathcal{G}]$ in $w\left(L^{1},L^{\infty}\right)$. ∎
Corollary 1.
The set $L^{1}(\mathcal{G})$ is $w\left(L^{1},L^{\infty}\right)$-closed in $L^{1}$.
Proof.
Let $X_{\lambda}\in L^{1}(\mathcal{G})$ and $X_{\lambda}\to X$ in $w\left(L^{1},L^{\infty}\right)$. This also implies that $X_{\lambda}\to\mathbb{E}[X|\mathcal{G}]$ in $w\left(L^{1},L^{\infty}\right)$ by Proposition 16. Since $w\left(L^{1},L^{\infty}\right)$ is Hausdorff, $X=\mathbb{E}[X|\mathcal{G}]\implies X\in L^{1}(\mathcal{G})$. Therefore $L^{1}(\mathcal{G})$ is $w\left(L^{1},L^{\infty}\right)$-closed in $L^{1}$. ∎
Proof of Proposition 6.
$(\mathrm{i})\Rightarrow(\mathrm{ii})$ and $(\mathrm{ii})\Rightarrow(\mathrm{iii})$: These are clear. $(\mathrm{iii})\Rightarrow(\mathrm{iv})$: We have
$\rho(X)=\rho(\mathbb{E}[X|\mathcal{G}])=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[D_{\mu}\mathbb{E}[X|\mathcal{G}]]=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[\mathbb{E}[D_{\mu}|\mathcal{G}]X].$
Therefore the set $\{\mathbb{E}[D_{\mu}|\mathcal{G}]:\mu\in\mathcal{S}_{\rho}\}\subseteq\widehat{\mathcal{S}}_{\rho}$ is $w(L^{1},L^{\infty})$-dense by Proposition 10. Since $L^{1}(\mathcal{G})$ is $w(L^{1},L^{\infty})$-closed by Corollary 1, we have $\widehat{\mathcal{S}}_{\rho}\subseteq L^{1}(\mathcal{G})$. $(\mathrm{iv})\Rightarrow(\mathrm{i})$: Define the map
$\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}:X\mapsto\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[D_{\mu}X].$
The mapping $\tilde{\rho}$ is clearly coherent and Fatou. By the tower property of conditional expectation, we have
$\rho(X)=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[D_{\mu}X]=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[D_{\mu}\mathbb{E}[X|\mathcal{G}]]=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}]).$
The final statement of the proposition follows by definition. ∎
Proof of Proposition 7.
Define $\tilde{\rho}:L^{\infty}\to\mathbb{R}$ by $\tilde{\rho}(X)=\gamma(X_{\#}\mathbb{P})$, $X\in L^{\infty}$. By definition, $\tilde{\rho}$ is coherent and Fatou. For $\mu\in\mathcal{S}$, define $\rho^{\mu}:L^{\infty}\to\mathbb{R}$ by $\rho^{\mu}(X)=\gamma(X_{\#}\mu)$, $X\in L^{\infty}$. It is straightforward to see that for $X,Y\in L^{\infty}$, $\rho^{\mu}(X)=\tilde{\rho}(Y)$ if $F_{\mu,X}=F_{\mathbb{P},Y}$. We will first prove the coherence of $\rho^{\mu}$. Translation invariance and positive homogeneity of $\rho^{\mu}$ is apparent. Let $X,Y\in L^{\infty}$ with $X\leq Y$, we have that $F_{\mu,Y}(x)\leq F_{\mu,X}(x)$ for all $x\in\mathbb{R}$ or $F_{\mu,X}^{-1}(u)\leq F_{\mu,Y}^{-1}(u)$ for all $u\in[0,1]$. Since $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, there exists a uniform $(0,1)$ random variable $U$. Define $X_{\mu}=F_{\mu,X}^{-1}(U)$ and $Y_{\mu}=F_{\mu,Y}^{-1}(U).$ We have $F_{\mu,X}=F_{\mathbb{P},X_{\mu}}$, $F_{\mu,Y}=F_{\mathbb{P},Y_{\mu}}$ and $X_{\mu}\leq Y_{\mu}.$ Thus, $\rho^{\mu}(X)=\tilde{\rho}(X_{\mu})\leq\tilde{\rho}(Y_{\mu})=\rho^{\mu}(Y).$ Therefore $\rho^{\mu}$ is monontonic.
Let $X,Y\in L^{\infty}$, by Dudley (2002, Theorem 11.7.5), there exists a Borel-measurable $f:(0,1)\to\mathbb{R}^{2}$ such that $f_{\#}\lambda=(X,Y)_{\#}\mu$, where $\lambda$ is the Lebesgue measure. Since $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, there exists a uniform $(0,1)$ random variable $U\in L^{\infty}$. Define $(X_{\mu},Y_{\mu})=f\circ U$, we have $(X,Y)_{\#}\mu=(X_{\mu},Y_{\mu})_{\#}\mathbb{P}.$ Therefore $F_{\mu,X}=F_{\mathbb{P},X_{\mu}}$, $F_{\mu,Y}=F_{\mathbb{P},Y_{\mu}}$ and $F_{\mu,X+Y}=F_{\mathbb{P},X_{\mu}+Y_{\mu}}$. Thus,
$\rho^{\mu}(X+Y)=\tilde{\rho}(X_{\mu}+Y_{\mu})\leq\tilde{\rho}(X_{\mu})+\tilde{\rho}(Y_{\mu})=\rho^{\mu}(X)+\rho^{\mu}(Y).$
Therefore, $\rho^{\mu}$ is subadditive.
Let $(X_{n})_{n\in\mathbb{N}}$ in $L^{\infty}$ be uniformly bounded and converge $\mathbb{P}$-a.s. to $X\in L^{\infty}$. Therefore, $(X_{n})_{n\in\mathbb{N}}$ converges $\mu$-a.s. to $X$. Thus,
$\lim_{n\to\infty}F^{-1}_{\mu,X_{n}}(q)=F^{-1}_{\mu,X}(q)$
for all the continuity points of $F^{-1}_{X,\mu}$. Since $(\Omega,\mathcal{F},\mathbb{P})$ is atomless, there exists a uniform $(0,1)$ random variable $U$. Define $X_{\mu}^{n}=F_{\mu,X_{n}}^{-1}(U)$, $X_{\mu}=F_{\mu,X}^{-1}(U)$. We have $F_{\mu,X_{n}}=F_{\mathbb{P},X_{\mu}^{n}}$, $F_{\mu,X}=F_{\mathbb{P},X_{\mu}}$. Since the set of discontinuity points of $F^{-1}_{X,\mu}$ is countable, $X_{\mu}^{n}$ converges $\mathbb{P}$-a.s. to $X_{\mu}$. Thus,
$\rho^{\mu}(X)=\tilde{\rho}(X_{\mu})\leq\liminf_{n\to\infty}\tilde{\rho}(X_{\mu}^{n})=\liminf_{n\to\infty}\rho^{\mu}(X_{n}).$
Thus proving Fatou continuity.
Finally, since
$\rho(X)=\sup_{\mu\in\mathcal{S}}\gamma(X_{\#}\mu)=\sup_{\mu\in\mathcal{S}}\rho^{\mu}(X),$
it follows that $\rho$ is the supremum of coherent and Fatou risk measures. Thus, $\rho$ is coherent and Fatou. The fact that $\rho$ is $\mathcal{G}$-law invariant is a straightforward consequence of the fact that $\mathcal{S}\subseteq\mathcal{E}(\mathbb{P}_{0}).$
Proof of the claims in Example 4.
Claim 1: The first equality is straightforward. To show the second equality, let $X\in L^{\infty}(\mathcal{H})$, from the dual representation of ES in Example 9, we get
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X]\leq\sup_{\mu\in\mathcal{S}_{\beta,\mathbb{P}}}\mathbb{E}^{\mu}[X]=\mathrm{ES}_{\beta}(X).$
For the other direction, let $A\in\mathcal{F}$ be an event independent of $\mathcal{G}$ satisfying $\mathbb{E}[X|A]=\mathrm{ES}_{\beta}(X)$ and $\mathbb{P}(A)=1-\beta$. The existence of such an event (called an $\beta$-tail event) is justified by Wang and Ziegel (2021, Lemmas A.4 and A.7). Let $\nu\in\mathcal{M}_{1}$ be given by $D_{\nu}=\mathds{1}_{A}/(1-\beta)$. Since $\mathds{1}_{A}$ is independent of $\mathcal{G}$, we have $\nu\in\mathcal{E}(\mathbb{P})$, and therefore $\nu\in\mathcal{S}$, and $\rho(X)\geq\mathbb{E}^{\nu}[X]=\mathbb{E}[X|A]=\mathrm{ES}_{\beta}(X)$. Claim 2: Take $X_{1}\in L^{\infty}(\mathcal{G})$, $X_{2}\in L^{\infty}(\mathcal{H})$, $X_{3}\in L^{\infty}$ with $X_{1}\overset{\mathrm{d}}{=}_{\mathbb{P}}X_{2}\overset{\mathrm{d}}{=}_{\mathbb{P}}X_{3}$. We have
$\rho(X_{3})$ $=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X_{3}]\geq\mathbb{E}[X_{3}]=\mathbb{E}[X_{1}]=\rho(X_{1});$
$\rho(X_{3})$ $=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X_{3}]\leq\sup_{\mu\in\mathcal{S}_{\beta,\mathbb{P}}}\mathbb{E}^{\mu}[X_{3}]=\mathrm{ES}_{\beta}[X_{3}]=\mathrm{ES}_{\beta}[X_{2}]=\rho(X_{2}),$
as desired. ∎
Lemma 3.
Föllmer and Schied (2016, Lemma A.32) Let $\mathcal{J}\subseteq\mathcal{F}$ be a sub-$\sigma$-algebra such that $(\Omega,\mathcal{J},\mathbb{P})$ is atomless. For all $X\in L^{1}(\mathcal{J})$ there exists a uniform $(0,1)$ random variable $U\in L^{\infty}(\mathcal{J})$ such that $X=F_{\mathbb{P},X}^{-1}(U)$  $\mathbb{P}$-a.s.
Proof of claim in Example 5.
This proof follows a very similar line of reasoning as Liu et al. (2022, Proposition 2); we reproduce the proof here as there are some nuances in this setting that need to be considered.
The $\beta$-tail measure $\widehat{\rho}^{\beta}:L^{\infty}\to\mathbb{R}$ generated by $\widehat{\rho}$ is defined by
$\widehat{\rho}^{\beta}(X)=\widehat{\rho}(F_{\mathbb{P},X}^{-1}\left(\beta+(1-\beta)U)\right),$
where $U$ is a uniform $(0,1)$ random variable. Since $\widehat{\rho}$ is law-invariant, this mapping is clearly well-defined. For $\mu\in\mathcal{M}_{1}$, define $\rho^{\mu}:L^{\infty}\to\mathbb{R}$ by $\rho^{\mu}(X)=\gamma(X_{\#}\mu)$, $X\in L^{\infty}$. It is straightforward to see that for $X,Y\in L^{\infty}$, $\rho^{\mu}(X)=\widehat{\rho}(Y)$ if $F_{\mu,X}=F_{\mathbb{P},Y}$. Recall that $\mathcal{S}_{\beta,\mathbb{P}}=\{\nu\in\mathcal{M}_{1}\mid({1-\alpha})D_{\nu}\leq 1\}$. By Liu et al. (2022, Proposition 2), we have
$\rho(X)\leq\sup_{\mu\in\mathcal{S}_{\beta,\mathbb{P}}}\rho^{\mu}(X)=\widehat{\rho}^{\beta}(X),$
[…]
$D_{\mu}=\frac{1}{1-\beta}\mathds{1}_{\{U_{X}>\beta\}}.$
Since $D_{\mu}$ is independent of $\mathcal{G}$, $\mathbb{E}[D_{\mu}|\mathcal{G}]=1$ and $\mu\in\mathcal{S}$. Let $Z=F_{\mathbb{P},X}^{-1}(\beta+(1-\beta)U_{X})$ and $x\in\mathbb{R}$, we have
$\mu(X>x)$ $=\mu(X>x,U_{X}>\beta)=\frac{1}{1-\beta}\mathbb{P}(X>x,U_{X}>\beta)=\min\left\{\frac{1}{1-\beta}\mathbb{P}({X>x}),1\right\},$
$F_{\mu,X}(x)$ $=\left(\frac{1}{1-\beta}(F_{\mathbb{P},X}(x)-\beta)\right)^{+}=F_{\mathbb{P},Z}(x).$
Implying,
$\widehat{\rho}^{\beta}(X)=\widehat{\rho}(Z)=\rho^{\mu}(X)\leq\rho(X),$
and $\rho(X)=\widehat{\rho}^{\beta}(X).$ Let $X\in L^{\infty}(\mathcal{G})$ and $\mu\in\mathcal{S}$, since $\mu\in\mathcal{E}(\mathbb{P})$, $F_{\mu,X}=F_{\mathbb{P},X}$. Therefore, $\rho^{\mu}(X)=\widehat{\rho}(X)$ and $\rho(X)=\widehat{\rho}(X)$. ∎
Appendix D An additional example accompanying Section 4
Example 10.
Let $\mathfrak{G}\subseteq\Sigma(\mathcal{F})$ be a collection of pair-wise independent $\sigma$-algebras. For each $\mathcal{G}\in\mathfrak{G}$, let $\rho_{\mathcal{G}}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ be a Fatou law-invariant coherent risk measure. Define the risk measure
$\rho(X)=\sup_{\mathcal{G}\in\mathfrak{G}}\rho_{\mathcal{G}}(\mathbb{E}[X|\mathcal{G}]),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Using a similar argument as Example 3 in Section 3.2, one can show that $\rho$ is a Fatou $\mathfrak{G}$-law-invariant coherent risk measure. To see how this example relates to Theorem 2, for each $\mathcal{G}\in\mathfrak{G}$, let $\iota_{\mathcal{G}}:\mathcal{M}_{1}(\mathcal{G})\to\mathcal{M}_{1}$ be the embedding given by $[\iota_{\mathcal{G}}\mu](A)=\mathbb{E}^{\mu}[\mathds{1}_{A}]$, $A\in\mathcal{F}$. Also, denote the supporting set of $\rho_{\mathcal{G}}$ by $\mathcal{S}_{\mathcal{G}}$. Define $\mathcal{S}=\operatorname*{conv}\left(\bigcup_{\mathcal{G}\in\mathfrak{G}}\iota_{\mathcal{G}}(\mathcal{S}_{\mathcal{G}})\right).$ By Proposition 6, $\sup_{\mu\in\iota_{\mathcal{G}}(\mathcal{S}_{\mathcal{G}})}\mathbb{E}^{\mu}[X]=\rho_{\mathcal{G}}(\mathbb{E}[X|\mathcal{G}])$ for all $X\in L^{\infty}$. Thus, $\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X]$, $X\in L^{\infty}.$ Since $\mathfrak{G}$ is independent, for all $\mathcal{G}\in\mathfrak{G}$, we have $\mathcal{S}^{\mathcal{G}}=\mathcal{S}_{\mathcal{G}}$, which is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closed and $\mathcal{G}$-law invariant. Therefore, the set $\mathcal{S}$ is a valid candidate for the set found in Theorem 2.
Appendix E Additional results and proofs accompanying Section 5
E.1 Additional results and proofs accompanying Section 5.1
Proof of Proposition 8.
(i)$\Rightarrow$(ii): For $X\in L^{\infty}$ define $\tau(X)=\rho(X)-\mathbb{E}\left[D_{\mu}X\right]$. The fact that $\tau$ is a $\mu$-coherent adjustment is easily verified. To check that $\tau$ is Fatou, let $(X_{n})_{n=1}^{\infty}$ be a uniformly bounded sequence in $L^{\infty}$ which converges $\mathbb{P}$-a.s. to some $X\in L^{\infty}$. By the Dominated Convergence Theorem (DCT) we have $\lim_{n\to\infty}\mathbb{E}\left[D_{\mu}X_{n}\right]=\mathbb{E}\left[D_{\mu}X\right]$. Therefore,
$\liminf_{n\to\infty}\tau(X_{n})=\liminf_{n\to\infty}(\rho(X_{n})-\mathbb{E}\left[D_{\mu}X_{n}\right])$ $\geq\liminf_{n\to\infty}\rho(X_{n})-\lim_{n\to\infty}\mathbb{E}\left[D_{\mu}X_{n}\right]$
$\geq\rho(X)-\mathbb{E}\left[D_{\mu}X\right]=\tau(X).$
The claim follows since $\rho(X)=\mathbb{E}\left[D_{\mu}X\right]+\tau(X)$ for all $X\in L^{\infty}$.
(ii)$\Rightarrow$(i): The facts that $\rho$ is a coherent risk measure and that its restriction is $\mathbb{E}^{\mu}$ are easily verified. Fatou continuity follows an argument similar to the one given above. ∎
We provide some facts regarding the representation of $\mu$-coherent adjustments.
Proposition 17.
Let $\tau:L^{\infty}\to\mathbb{R}$ and $\mu\in\mathcal{M}_{1}(\mathcal{G})$. The following are equivalent.
  1. (i)
    The mapping $\tau\in\mathrm{CA}(\mu)$;
  2. (ii)
    there exists a $w(\mathcal{M},L^{\infty})$-closed and convex set $\mathcal{S}_{\tau}\subseteq\mathcal{E}(\mu)$ such that
    $\tau(X)=\sup_{\nu\in\mathcal{S}_{\tau}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (18)
  3. (iii)
    there exists $\mathcal{R}\subseteq\mathcal{E}(\mu)$ such that $\tau(X)=\sup_{\nu\in\mathcal{R}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\}$, $X\in L^{\infty}$.
In case (ii) holds, the closed and convex set $\mathcal{S}_{\tau}$ is uniquely given by
$\left\{\nu\in\mathcal{M}_{1}:\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\leq\tau(X)\text{ for all }X\in L^{\infty}\right\}.$
Proof.
$(\mathrm{i})\Rightarrow(\mathrm{ii})$: Assume that $\tau\in\mathrm{CA}(\mu)$, by Proposition 8, the mapping $\rho=\mathbb{E}^{\mu}+\tau$ is a Fatou coherent risk measure. We have
$\mathbb{E}^{\mu}\left[X\right]+\tau(X)=\sup_{\nu\in\mathcal{S}_{\rho}}\mathbb{E}^{\nu}\left[X\right],\hskip 3.0ptX\in L^{\infty}.$
By Proposition 13, $\mathcal{S}^{\mathcal{G}}_{\rho}$ must be contained in the supporting set of $\rho|_{L^{\infty}(\mathcal{G})}=\mathbb{E}^{\mu}|_{L^{\infty}(\mathcal{G})}$, and therefore $\mathcal{S}^{\mathcal{G}}_{\rho}=\{\mu\}$ (meaning $\mathcal{S}_{\rho}\subseteq\mathcal{E}(\mu)$). Note that
$\mathbb{E}^{\mu}[X]+\tau(X)=\mathbb{E}^{\mu}[X]+\sup_{\nu\in\mathcal{S}_{\rho}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
implying $\tau(X)=\sup_{\nu\in\mathcal{S}_{\rho}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\}$, $X\in L^{\infty}$. Therefore (18) holds with $\mathcal{S}_{\tau}=\mathcal{S}_{\rho}$. Furthermore,
$\mathcal{S}_{\tau}=\mathcal{S}_{\rho}$ =$=\left\{\nu\in\mathcal{M}_{1}\mid\mathbb{E}^{\nu}\left[X\right]\leq\rho(X)\text{ for all }X\in L^{\infty}\right\}$
=$=\left\{\nu\in\mathcal{M}_{1}\mid\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\leq\tau(X)\text{ for all }X\in L^{\infty}\right\}.$
The uniqueness follows directly from Proposition 10.
$(\mathrm{ii})\Rightarrow(\mathrm{iii})$: Obvious. $(\mathrm{iii})\Rightarrow(\mathrm{i})$: Let $\mathcal{R}\subseteq\mathcal{M}_{1}$ have the stated properties. Proving $\tau$ is a $\mu$-coherent adjustment is straightforward. Let $(X_{n})_{n=1}^{\infty}$ be a bounded sequence that converges $\mathbb{P}$-a.s. to $X$. By the DCT, $\lim_{n\to\infty}\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X_{n}\right]=\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X\right]$. Therefore,
$\tau(X)=\sup_{\nu\in\mathcal{R}}\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X\right]=\sup_{\nu\in\mathcal{R}}\liminf\limits_{n\to\infty}\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X_{n}\right]$ $\leq\liminf\limits_{n\to\infty}\sup_{\nu\in\mathcal{R}}\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X_{n}\right]$
$=\liminf\limits_{n\to\infty}\tau(X_{n}),$
proving the Fatou continuity of $\tau$. ∎
Corollary 2.
Let $\mu\in\mathcal{M}_{1}(\mathcal{G})$ and $\tau:L^{\infty}\to\mathbb{R}$ be a $\mu$-coherent adjustment. The following are equivalent.
  1. (i)
    The mapping $\tau\in\mathrm{CA}(\mu)$;
  2. (ii)
    the mapping $\tau$ is $w(L^{\infty},\mathcal{M})$-lower semicontinuous;
  3. (iii)
    the mapping $\mathbb{E}^{\mu}+\tau$ is Fatou;
  4. (iv)
    the mapping $\mathbb{E}^{\mu}+\tau$ is $w(L^{\infty},\mathcal{M})$-lower semicontinuous.
Proof.
$(\mathrm{i})\Rightarrow(\mathrm{ii})$: This follows from the representation given in Proposition 17 and by the fact that the supremum of continuous functions is lower semicontinuous. $(\mathrm{ii})\Rightarrow(\mathrm{iv})$: The sum of lower semicontinuous functions is lower semicontinuous. $(\mathrm{iv})\Rightarrow(\mathrm{iii})$: This follows since $\mathbb{E}^{\mu}+\tau$ is a coherent risk measure that is lower $w(L^{\infty},\mathcal{M})$-lower semicontinuous. $(\mathrm{iii})\Rightarrow(\mathrm{i})$: Since $\mathbb{E}^{\mu}+\tau$ is Fatou, we can use a similar argument as Proposition 17 to show that $\tau$ has the representation (18). Thus by Proposition 17, the mapping $\tau$ is Fatou. ∎
Proof of Theorem 3.
Let $\rho$ be a Fatou coherent risk measure. For $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$, we define $\mathcal{S}(\mu)=\mathcal{E}(\mu)\cap\mathcal{S}_{\rho}$, note that $\mathcal{S}_{\rho}=\bigcup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\mathcal{S}(\mu)$. We have
$\rho(X)=\sup_{\nu\in\mathcal{S}_{\rho}}\mathbb{E}\left[D_{\mu}X\right]$ $=\sup_{\nu\in\mathcal{S}_{\rho}}\left\{\mathbb{E}\left[\mathbb{E}[D_{\nu}|\mathcal{G}]X\right]+\mathbb{E}[(D_{\nu}-\mathbb{E}[D_{\nu}|\mathcal{G}])X]\right\}$
$=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\sup_{\nu\in\mathcal{S}(\mu)}\left\{\mathbb{E}\left[D_{\mu}X\right]+\mathbb{E}\left[(D_{\nu}-D_{\mu})X\right]\right\}$
$=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\}$, $X\in L^{\infty}$. The fact that $\tau_{\mu}\in\mathrm{CA}(\mu)$ follows from Proposition 17. It is straightforward to see that $\mathcal{S}_{\rho}^{\mathcal{G}}$ is convex. For the converse, note that (12) is the supremum of coherent risk measures (by Proposition 8; see (11)) and thus a coherent risk measure. By Corollary 2, $\mathbb{E}^{\mu}+\tau_{\mu}$ is $w(L^{\infty},\mathcal{M})$-lower semicontinuous for all $\mu\in\mathcal{R}$. As the supremum of $w(L^{\infty},\mathcal{M})$-lower semicontinuous and coherent risk measures, $\rho$ is also one, and hence it is Fatou. Let $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$. We have that $\mathcal{R}$ in (12) is dense in $\mathcal{S}_{\tilde{\rho}}$, and the last claim now follows from the same argument in the proof of Theorem 1. ∎
E.2 Additivity on the subspace
In this section, we discuss coherent risk measures that are not only subadditive but also additive on $L^{\infty}(\mathcal{G})$, which turn out to have a strong connection to coherent adjustments.
For $\rho:L^{\infty}\to\mathbb{R}$, we say that $\rho$ is $\mathcal{G}$-additive if $\rho(X+Y)=\rho(X)+\rho(Y)$ for all $X,Y\in L^{\infty}(\mathcal{G})$. Additivity has the interpretation as risk neutrality under a particular probability, an idea that goes back to at least de Finetti (1931); see Wakker (2010, p. 23) for discussions on this. Hence, $\mathcal{G}$-additivity reflects that the risk measure is risk neutral on the subset $L^{\infty}(\mathcal{G})$ of $L^{\infty}$. In other words, such a risk measure is neutral for risks without uncertainty; recall that in the context of Section 2.2, random losses in $L^{\infty}(\mathcal{G})$ do not have uncertainty in their distributions.
For $X\in L^{\infty}$, we say that $\rho$ is linear along $X$ if $\rho(\lambda X)=\lambda\rho(X)$ for all $\lambda\in\mathbb{R}$; see Bellini et al. (2021) and Liebrich and Munari (2022) for more discussions on this technical property. The following proposition gives necessary and sufficient conditions for dropping the supremum in representation (12) in the case of $\mathcal{G}$-law invariance.
Proposition 18.
Let $\rho:L^{\infty}\to\mathbb{R}$ be a Fatou coherent risk measure. The following are equivalent.
  1. (i)
    The mapping $\rho$ is $\mathcal{G}$-additive and $\mathcal{G}$-law invariant;
  2. (ii)
    the mapping $\rho$ is linear along a nonconstant $X\in L^{\infty}(\mathcal{G})$ and $\mathcal{G}$-law invariant;
  3. (iii)
    there exists $\tau\in\mathrm{CA}(\mathbb{P}|_{\mathcal{G}})$ such that
    $\rho(X)=\mathbb{E}[X]+\tau(X),\hskip 3.0ptX\in L^{\infty}.$
Proof.
(i)$\Rightarrow$(ii): Since $\rho$ is $\mathcal{G}$-additive and coherent, it is linear on $L^{\infty}(\mathcal{G})$. Since $L^{\infty}(\mathcal{G})$ contains non-constant random variables, (ii) holds. (ii)$\Rightarrow$(iii): The fact that $\rho|_{L^{\infty}(\mathcal{G})}=\mathbb{E}|_{L^{\infty}(\mathcal{G})}$ follows from Theorem 4.5 and Theorem 4.7 of Bellini et al. (2021). Using Proposition 8, we prove the claim. (iii)$\Rightarrow$(i): This is clear by definition. ∎
E.3 Uniqueness of the representation in Theorem 3
A natural question would be, under what conditions is representation (12) unique? When the coherent risk measure is Lebesgue continuous, there is a definite answer to this. We first need some results before we can state and prove the theorem regarding uniqueness.
If $\left(\mathfrak{X},\mathfrak{Y},\langle\;,\;\rangle\right)$ is a dual pairing, we say that $\mathcal{R}\subseteq\mathfrak{X}$ is $w(\mathfrak{X},\mathfrak{Y})$-precompact if the $w(\mathfrak{X},\mathfrak{Y})$-closure of $\mathcal{R}$ is $w(\mathfrak{X},\mathfrak{Y})$-compact.
The next proposition, parallel to Proposition 17, characterizes Lebesgue-continuous $\mu$-coherent adjustments.
Proposition 19.
Let $\tau:L^{\infty}\to\mathbb{R}$ and $\mu\in\mathcal{M}_{1}(\mathcal{G})$. The following are equivalent.
  1. (i)
    The mapping $\tau$ is a Lebesgue-continuous $\mu$-coherent adjustment;
  2. (ii)
    there exists a $w(\mathcal{M},L^{\infty})$-compact and convex set $\mathcal{S}_{\tau}\subseteq\mathcal{E}(\mu)$ such that
    $\tau(X)=\sup_{\nu\in\mathcal{S}_{\tau}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (19)
  3. (iii)
    there exists $w(\mathcal{M},L^{\infty})$-precompact $\mathcal{R}\subseteq\mathcal{E}(\mu)$ such that
    $\tau(X)=\sup_{\nu\in\mathcal{R}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\}\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
In case (ii) holds, the set $\mathcal{S}_{\tau}$ is uniquely given by $\left\{\nu\in\mathcal{M}_{1}:\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\leq\tau(X)\text{ for all }X\in L^{\infty}\right\}$.
Proof.
Everything follows the same line of reasoning as Proposition 17, except for proving Lebesgue continuity in $(\mathrm{iii})\Rightarrow(\mathrm{i})$. Let $(X_{n})_{n=1}^{\infty}$ be a sequence in $L^{\infty}$ of uniformly bounded random variables that converge $\mathbb{P}$-a.s. to $X\in L^{\infty}$. Since $X\mapsto\sup_{\nu\in\mathcal{R}}\mathbb{E}[D_{\nu}X]$ is a Lebesgue-continuous coherent risk measure, we have
$\lim_{n\to\infty}\tau(X_{n})=\lim_{n\to\infty}\sup_{\nu\in\mathcal{R}}\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X_{n}\right]$ $=\lim_{n\to\infty}\sup_{\nu\in\mathcal{R}}\mathbb{E}\left[D_{\nu}X_{n}\right]-\lim_{n\to\infty}\mathbb{E}[D_{\mu}X_{n}]$
$=\sup_{\nu\in\mathcal{R}}\mathbb{E}[D_{\nu}X]-\mathbb{E}[D_{\mu}X]=\tau(X).$
Proving the Lebesgue continuity of $\tau$. ∎
Let $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ be $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closed and convex, and $(\tau_{\mu})_{\mu\in\mathcal{R}}$ be a CA-assignment. We say that $(\tau_{\mu})_{\mu\in\mathcal{R}}$ is Lebesgue continuous if $\tau_{\mu}$ is Lebesgue continuous for all $\mu\in\mathcal{R}$. We say that $(\tau_{\mu})_{\mu\in\mathcal{R}}$ is regular if
  1. (a)
    For any fixed $X\in L^{\infty}$, concavity of $\mu\mapsto\tau_{\mu}(X)$ holds; that is, for all $\mu_{1},\mu_{2}\in\mathcal{R}$ and $\lambda\in(0,1)$ we have
    $\lambda\tau_{\mu_{1}}(X)+(1-\lambda)\tau_{\mu_{2}}(X)\leq\tau_{\lambda\mu_{1}+(1-\lambda)\mu_{2}}(X).$
    See e.g., Wang et al. (2020) for concavity for distributional mixtures in the context of Choquet integrals.
  2. (b)
    A weak form of semicontinuity holds: If $(\mu_{\lambda})_{\lambda\in\Lambda}$ is a net in $\mathcal{R}$ that converges to $\mu\in\mathcal{R}$ in $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$, then for all $X\in L^{\infty}$,
    $\tau_{\mu}(X)\geq\liminf_{\lambda}\tau_{\mu_{\lambda}}(X).$
See Bredon (1993) for background information on nets, subnets, and their convergence.
Proposition 20.
Let $\rho:L^{\infty}\to\mathbb{R}$ be a Lebesgue-continuous coherent risk measure. There exists a unique $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-compact and convex set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a unique regular Lebesgue-continuous CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ such that $\rho$ can be represented by equation (12).
Proof.
Let $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$, by Proposition 14, we have that $\mathcal{S}^{\mathcal{G}}_{\rho}=\mathcal{S}_{\tilde{\rho}}$ is compact. Let $(\tau_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ be the same as in the construction of representation (12) in the proof of Theorem 3. We claim that for all $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$, $\mathcal{S}(\mu)$ is $w(\mathcal{M},L^{\infty})$-compact and convex. Fix some $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$, convexity of $\mathcal{S}(\mu)$ follows from the convexity of $\mathcal{S}_{\rho}$ and the convexity of $\mathcal{E}(\mu)$. For compactness note that $\mathcal{S}(\mu)$ is $w(\mathcal{M},L^{\infty})$-precompact as $\mathcal{S}(\mu)\subseteq\mathcal{S}_{\rho}$. Therefore, compactness of $\mathcal{S}(\mu)$ would follow if we prove that $\mathcal{S}(\mu)$ is $w(\mathcal{M},L^{\infty})$-closed. This can be seen from the fact that $\mathcal{S}_{\rho}$ and $\mathcal{E}(\mu)$ are both $w(\mathcal{M},L^{\infty})$-closed. From the construction in 3, $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\}$, $X\in L^{\infty}$, thus, by Proposition 19, $\tau_{\mu}$ is Lebesgue continuous.
We will now show that $(\tau_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ is regular. First, to show concavity (a), let $\mu_{1},\mu_{2}\in\mathcal{S}_{\rho}^{\mathcal{G}}$, $\nu_{1}\in\mathcal{S}(\mu_{1}),\nu_{2}\in\mathcal{S}(\mu_{2})$, $X\in L^{\infty}$, and $\lambda\in(0,1)$. Define $\upsilon=\lambda\mu_{1}+(1-\lambda)\mu_{2}$, since $\mathcal{S}_{\rho}$ is convex we have $\lambda\nu_{1}+(1-\lambda)\nu_{2}\in\mathcal{S}(\upsilon)$, and thus
$\mathbb{E}\left[\left(D_{\nu_{1}}-D_{\mu_{1}}\right)\lambda X\right]+\mathbb{E}\left[\left(D_{\nu_{2}}-D_{\mu_{2}}\right)(1-\lambda)X\right]\leq\tau_{\upsilon}(X).$ (20)
Taking the supremum of the left side of (20) over $(\nu_{1},\nu_{2})\in\mathcal{S}(\mu_{1})\times\mathcal{S}(\mu_{2})$ obtains the result. To show semicontinuity (b), let $(\mu_{\lambda})_{\lambda\in\Lambda}$ be a net in $\mathcal{S}_{\rho}^{\mathcal{G}}$ and $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$ such that $\mu_{\lambda}\to\mu$ in $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$ and let $X\in L^{\infty}$. For each $\lambda$ choose $\nu_{\lambda}^{*}\in\mathcal{S}(\mu_{\lambda})$ such that $\tau_{\mu_{\lambda}}(X)=\mathbb{E}[(D_{\nu_{\lambda}^{*}}-D_{\mu_{\lambda}})X]$, which is possible since $\mathcal{S}(\mu_{\lambda})$ is $w(\mathcal{M},L^{\infty})$-compact. Since the net $(\nu_{\lambda}^{*})_{\lambda\in\Lambda}$ takes values in the $w(\mathcal{M},L^{\infty})$-compact set $\mathcal{S}_{\rho}$, there exists a convergent subnet $(\nu_{\gamma}^{*})_{\gamma\in\Gamma}$. Denote the limit of this subnet by $\nu\in\mathcal{S}_{\rho}$. For $Y\in L^{\infty}(\mathcal{G})$, by using the tower property and the convergence, we get
$\mathbb{E}\left[\mathbb{E}[D_{\nu}|\mathcal{G}]Y\right]=\mathbb{E}\left[D_{\nu}Y\right]=\lim_{\gamma}\mathbb{E}\left[D_{\nu_{\gamma}^{*}}Y\right]=\lim_{\gamma}\mathbb{E}\left[D_{\mu_{\gamma}}Y\right]=\mathbb{E}[D_{\mu}Y].$
Thus $\nu\in\mathcal{S}(\mu)$ and
$\tau_{\mu}(X)\geq\mathbb{E}\left[(D_{\nu}-D_{\mu})X\right]=\lim_{\gamma}\mathbb{E}\left[\left(D_{\nu^{*}_{\gamma}}-D_{\mu_{\gamma}}\right)X\right]$ $=\lim_{\gamma}\tau_{\mu_{\gamma}}(X)$
$=\liminf_{\gamma}\tau_{\mu_{\gamma}}(X)\geq\liminf_{\lambda}\tau_{\mu_{\lambda}}(X).$
This shows that $(\tau_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ is regular.
Now we prove the uniqueness of $(\tau_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ by assuming that there exists another regular Lebesgue-continuous CA-assignment $(\tilde{\tau}_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{S}^{\mathcal{G}}_{\rho}}\left\{\mathbb{E}^{\mu}[X]+\tilde{\tau}_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
We will show that $\tau_{\mu}=\tilde{\tau}_{\mu}$ for all $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$. For each $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$, let $\mathcal{R}(\mu)\subseteq\mathcal{M}_{1}$ be the $w(\mathcal{M},L^{\infty})$-compact convex set that gives representation (19) in Proposition 19 for $\tilde{\tau}_{\mu}$. Let $\tilde{\mathcal{S}}=\bigcup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\mathcal{R}(\mu)\subseteq\mathcal{M}_{1}$. Note that
$\sup_{\nu\in\tilde{\mathcal{S}}}\mathbb{E}^{\nu}[X]=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\left\{\sup_{\nu\in\mathcal{R}(\mu)}\mathbb{E}^{\nu}[X]\right\}$ $=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\left\{\mathbb{E}^{\mu}[X]+\sup_{\nu\in\mathcal{R}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\right\}\right\}$
$=\sup_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}\left\{\mathbb{E}^{\mu}[X]+\tilde{\tau}_{\mu}(X)\right\}=\rho(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Therefore $\operatorname*{conv}(\tilde{\mathcal{S}})$ is $w(\mathcal{M},L^{\infty})$-dense in $\mathcal{S}_{\rho}$ by Propostion 10. We will show that regularity of $(\tilde{\tau}_{\mu})_{\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}}$ implies that $\tilde{\mathcal{S}}$ is convex and $w(\mathcal{M},L^{\infty})$-closed. Let $\nu_{1}\in\mathcal{R}(\mu_{1})$ and $\nu_{2}\in\mathcal{R}(\mu_{2})$ and $\lambda\in(0,1)$. We have, for $X\in L^{\infty}$,
$\mathbb{E}\left[\left(\lambda D_{\nu_{1}}+(1-\lambda)D_{\nu_{2}}-\lambda D_{\mu_{1}}-(1-\lambda)D_{\mu_{2}}\right)X\right]$ $=\mathbb{E}\left[\left(D_{\nu_{1}}-D_{\mu_{1}}\right)\lambda X\right]+\mathbb{E}\left[\left(D_{\nu_{2}}-D_{\mu_{2}}\right)(1-\lambda)X\right]$
$\leq\tilde{\tau}_{\mu_{1}}(\lambda X)+\tilde{\tau}_{\mu_{2}}((1-\lambda)X)\leq\tau_{\lambda D_{\mu_{1}}+(1-\lambda)D_{\mu_{2}}}(X).$
Therefore, $\lambda\nu_{1}+(1-\lambda)\nu_{2}\in\mathcal{R}(\lambda\mu_{1}+(1-\lambda)\mu_{2})\subseteq\tilde{\mathcal{S}}$ by Proposition 19. This implies that $\tilde{\mathcal{S}}$ is convex. Let $(\nu_{\lambda})_{\lambda\in\Lambda}$ be a net taking values in $\tilde{\mathcal{S}}$ and $\nu\in\mathcal{M}_{1}$ such that $\lim_{\lambda}\nu_{\lambda}=\nu$ in $w(\mathcal{M},L^{\infty})$, and define $\mu_{\lambda}=\nu_{\lambda}|_{\mathcal{G}}$ and $\mu=\nu|_{\mathcal{G}}$. By Proposition 16, $\lim_{\lambda}\mu_{\lambda}=\mu$ in $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$. Since $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-closed, we have that $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$. For $X\in L^{\infty}$,
$\mathbb{E}\left[\left(D_{\nu}-D_{\mu}\right)X\right]=\lim_{\lambda}\mathbb{E}\left[\left(D_{\nu_{\lambda}}-D_{\mu_{\lambda}}\right)X\right]\leq\liminf_{\lambda}\tilde{\tau}_{\mu_{\lambda}}(X)\leq\tilde{\tau}_{\mu}(X),$
implying that $\nu\in\mathcal{R}(\mu)\subseteq\tilde{\mathcal{S}}$ and $\tilde{\mathcal{S}}$ is $w(\mathcal{M},L^{\infty})$-closed. Therefore $\tilde{\mathcal{S}}=\mathcal{S}_{\rho}$, which means $\mathcal{S}(\mu)=\mathcal{R}(\mu)$ for all $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$, implying $\tilde{\tau}_{\mu}=\tau_{\mu}$ for all $\mu\in\mathcal{S}_{\rho}^{\mathcal{G}}$. ∎
E.4 Additional results and proofs accompanying Section 5.2
Proof of Lemma 1.
The only if is clear. Let $Y_{1},Y_{2}\in L^{\infty}(\mathcal{G})$ with $Y_{1}\overset{\mathrm{d}}{=}_{\mathbb{P}}Y_{2}$ and $Z\in\ker(\mathcal{G})$. Define $X=Z+Y_{1}$, we have $\mathbb{E}[X|\mathcal{G}]=Y_{1}$ therefore
$\rho(Z+Y_{1})=\rho(X)=\rho(X-\mathbb{E}[X|\mathcal{G}]+Y_{2})=\rho(Z+Y_{2}).$
Therefore $\rho$ is strongly $\mathcal{G}$-law invariant. ∎
The proof of Theorem 4 relies on several results we present below.
Lemma 4.
Let $X\in L^{\infty}$ (resp. $L^{1}$), $Y\in L^{1}$ (resp. $L^{\infty}$) and $Z\in L^{1}(\mathcal{G})$ (resp. $L^{\infty}(\mathcal{G})$) with $Y\overset{\mathrm{d}}{=}_{\mathbb{P}}Z$. Then there exists $W\in L^{\infty}(\mathcal{G})$ (resp. $L^{1}(\mathcal{G})$) such that $X\overset{\mathrm{d}}{=}_{\mathbb{P}}W$ and $\mathbb{E}[XY]=\mathbb{E}[WZ].$
Proof.
[…]
For all $n\in\mathbb{N}$, $W_{p_{n}}\overset{\mathrm{d}}{=}_{\mathbb{P}}X$, implying that there exists an $M>0$ such that $\mathbb{P}(\sup_{n\in\mathbb{N}}|W_{p_{n}}|\leq M)=1$. Thus,
$\sup_{n\in\mathbb{N}}|W_{p_{n}}Z|\leq M|Z|\in L^{1}(\mathcal{G})$
and $(W_{p_{n}}Z)_{n=1}^{\infty}$ is uniformly integrable. All of this combines to show that $\lim_{n\to\infty}\mathbb{E}[W_{p_{n}}Z]=\mathbb{E}[W_{p}Z].$ Therefore, the function
$\zeta:[0,1]\to\mathbb{R}:p\mapsto\mathbb{E}[W_{p}Z]$
is continuous. Note that
$\zeta(1)=\mathbb{E}\left[F_{\mathbb{P},X}^{-1}(1-U_{Z})F_{\mathbb{P},Y}^{-1}(U_{Z})\right]\leq\mathbb{E}[XY]\leq\mathbb{E}\left[F_{\mathbb{P},X}^{-1}(U_{Z})F_{\mathbb{P},Y}^{-1}(U_{Z})\right]=\zeta(0).$
Thus, by the intermediate value theorem, there exists a $p_{0}\in[0,1]$ such that $\mathbb{E}[W_{p_{0}}Z]=\mathbb{E}[XY].$ Let $X\in L^{1}$, $Y\in L^{\infty}$ and $Z\in L^{\infty}(\mathcal{G})$ with $Y\overset{\mathrm{d}}{=}_{\mathbb{P}}Z$. The only thing that has to be checked is the uniform integrability condition. Let $M=\|Z\|_{\infty}+1$ and $K>0$, we have
$\mathbb{E}\left[|W_{p_{n}}Z|;|W_{p_{n}}Z|>K\right]\leq M\mathbb{E}\left[|W_{p_{n}}|;|W_{p_{n}}|>K/M\right]=M\mathbb{E}\left[|X|;|X|>K/M\right]$
for all $n\in\mathbb{N}$. This shows that $(W_{p_{n}}Z)_{n=1}^{\infty}$ is uniformly integrable. ∎
For a $\mathcal{G}$-law invariant set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$, we say that the CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ is $\mathcal{G}$-law invariant if for all $\mu,\nu\in\mathcal{R}$ with $D_{\mu}\overset{\mathrm{d}}{=}_{\mathbb{P}}D_{\nu}$, it holds that $\tau_{\mu}=\tau_{\nu}$.
Proposition 21.
Let $\rho$ be a Fatou coherent risk measure. The following are equivalent.
  1. (i)
    The risk measure $\rho$ is strongly $\mathcal{G}$-law invariant;
  2. (ii)
    the set $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $\mathcal{G}$-law invariant and for all $\mu\in\mathcal{S}_{\rho}$ and $Y\in L^{1}(\mathcal{G})$ with $\mathbb{E}[D_{\mu}|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$, it holds that
    $D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}]+Y\in\widehat{\mathcal{S}}_{\rho}.$
  3. (iii)
    There exists a convex $\mathcal{G}$-law invariant set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a $\mathcal{G}$-law invariant CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ such that
    $\rho(X)=\sup_{\mu\in\mathcal{R}}\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Proof.
$(\mathrm{i})\Rightarrow(\mathrm{ii}):$ We will prove that $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $\mathcal{G}$-law invariant after proving the latter claim. Let $\mu\in\mathcal{S}_{\rho}$ and let $Y\in L^{1}(\mathcal{G})$ with $\mathbb{E}[D_{\mu}|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$ such that
$D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}]+Y\notin\widehat{\mathcal{S}}_{\rho}.$
By the Hahn-Banach separation Theorem, there exists $X\in L^{\infty}$ such that
$\rho(X)<\mathbb{E}[(D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}]+Y)X].$
By Lemma 4, find $Z\in L^{\infty}(\mathcal{G})$ with $\mathbb{E}[X|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Z$ such that $\mathbb{E}[Y\mathbb{E}[X|\mathcal{G}]]=\mathbb{E}[\mathbb{E}[D_{\mu}|\mathcal{G}]Z]$. We have
$\mathbb{E}[(D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}]+Y)X]$ $=\mathbb{E}[(D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}])X]+\mathbb{E}[YX]$
$=\mathbb{E}[(D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}])X]+\mathbb{E}[Y\mathbb{E}[X|\mathcal{G}]]$
$=\mathbb{E}[D_{\mu}(X-\mathbb{E}[X|\mathcal{G}])]+\mathbb{E}[\mathbb{E}[D_{\mu}|\mathcal{G}]Z]$
$=\mathbb{E}[D_{\mu}(X-\mathbb{E}[X|\mathcal{G}])]+\mathbb{E}[D_{\mu}Z]$
$=\mathbb{E}[D_{\mu}(X-\mathbb{E}[X|\mathcal{G}]+Z)]$
$=\mathbb{E}^{\mu}[X-\mathbb{E}[X|\mathcal{G}]+Z].$
Therefore, $\rho(X)<\mathbb{E}^{\mu}[X-\mathbb{E}[X|\mathcal{G}]+Z]\leq\rho(X-\mathbb{E}[X|\mathcal{G}]+Z)$, implying that $\rho$ is not strongly $\mathcal{G}$-law invariant. We will now prove that $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $\mathcal{G}$-law invariant. Let $\mu\in\mathcal{S}_{\rho}$ and $Y\in L^{1}(\mathcal{G})$ such that $\mathbb{E}[D_{\mu}|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$. We have $D_{\mu}-\mathbb{E}[D_{\mu}]+Y\in\widehat{\mathcal{S}}_{\rho}$ and $Y=\mathbb{E}[D_{\mu}-\mathbb{E}[D_{\mu}|\mathcal{G}]+Y|\mathcal{G}]\in\widehat{\mathcal{S}}_{\rho}^{\mathcal{G}}$.
$(\mathrm{ii})\Rightarrow(\mathrm{iii}):$ Since $\rho$ is a Fatou coherent risk measure, by Theorem 3, we have
$\rho(X)=\sup_{\mu\in\mathcal{S}^{\mathcal{G}}_{\rho}}\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\}$, $X\in L^{\infty}$ and $\mathcal{S}(\mu)=\mathcal{E}(\mu)\cap\mathcal{S}_{\rho}$. The set $\mathcal{S}_{\rho}^{\mathcal{G}}$ is $\mathcal{G}$-law invariant by assumption and is convex as $\mathcal{S}_{\rho}$ is convex. Let $\mu_{1},\mu_{2}\in\mathcal{S}_{\rho}^{\mathcal{G}}$ with $D_{\mu_{1}}\overset{\mathrm{d}}{=}_{\mathbb{P}}D_{\mu_{2}}$ and $\nu_{1}\in\mathcal{S}(\mu_{1})$. By assumption, there exists $\upsilon\in\mathcal{S}_{\rho}$ with $D_{\upsilon}=D_{\nu_{1}}-D_{\mu_{1}}+D_{\mu_{2}}$. This implies that $\upsilon\in\mathcal{E}(\mu_{2})$ and thus $\upsilon\in\mathcal{S}(\mu_{2})$. Yet we have $D_{\upsilon}-D_{\mu_{2}}=D_{\nu_{1}}-D_{\mu_{1}}$, implying that for all $X\in L^{\infty}$,
$\tau_{\mu_{2}}(X)=\sup_{\upsilon\in\mathcal{S}(\mu_{2})}\mathbb{E}[(D_{\upsilon}-D_{\mu_{2}})X]\geq\sup_{\upsilon\in\mathcal{S}(\mu_{1})}\mathbb{E}[(D_{\upsilon}-D_{\mu_{1}})X]=\tau_{\mu_{1}}(X).$
By symmetry, we get the reverse inequality.
$(\mathrm{iii})\Rightarrow(\mathrm{i}):$ Let $\mu\in\mathcal{R}$, $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$ such that $\mathbb{E}[X|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$, define $W=X-\mathbb{E}[X|\mathcal{G}]+Y$. By Lemma 4, find $Z\in L^{1}(\mathcal{G})$ such $D_{\mu}\overset{\mathrm{d}}{=}_{\mathbb{P}}Z$ and $\mathbb{E}[D_{\mu}Y]=\mathbb{E}[Z\mathbb{E}[X|\mathcal{G}]]$. Let $\nu\in\mathcal{M}_{1}(\mathcal{G})$ be such that $D_{\nu}=Z$. Since $\mathcal{R}$ is $\mathcal{G}$-law invariant, $\nu\in\mathcal{R}$. We have
$\mathbb{E}^{\mu}[W]+\tau_{\mu}(W)=\mathbb{E}[D_{\mu}Y]+\tau_{\mu}(X)=\mathbb{E}[D_{\nu}\mathbb{E}[X|\mathcal{G}]]+\tau_{\nu}(X)=\mathbb{E}[D_{\nu}X]+\tau_{\nu}(X)=\mathbb{E}^{\nu}[X]+\tau_{\nu}(X).$
By Lemma 4, find $Z\in L^{1}(\mathcal{G})$ such $D_{\mu}\overset{\mathrm{d}}{=}_{\mathbb{P}}Z$ and $\mathbb{E}[D_{\mu}\mathbb{E}[X|\mathcal{G}]]=\mathbb{E}[ZY]$. Let $\nu\in\mathcal{M}_{1}(\mathcal{G})$ be such that $D_{\nu}=Z$. Since $\mathcal{R}$ is $\mathcal{G}$-law invariant, $\nu\in\mathcal{R}$. We have
$\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)=\mathbb{E}[D_{\mu}\mathbb{E}[X|\mathcal{G}]]+\tau_{\mu}(W)=\mathbb{E}[D_{\nu}Y]+\tau_{\nu}(W)$ $=\mathbb{E}[D_{\nu}W]+\tau_{\nu}(W)$
$=\mathbb{E}^{\nu}[W]+\tau_{\nu}(W).$
Therefore, $\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X):\mu\in\mathcal{R}\}=\{\mathbb{E}^{\mu}[W]+\tau_{\mu}(W):\mu\in\mathcal{R}\},$ and $\rho(X)=\rho(W).$
Proof of Theorem 4.
$(\mathrm{i})\Rightarrow(\mathrm{ii})$: By Proposition 21, we have that there exist a $\mathcal{G}$-law-invariant convex $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a $\mathcal{G}$-law-invariant CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
By Föllmer and Schied (2016, Theorem 4.62), for $Q\in\mathcal{M}_{B}$ and $\mu\in\mathcal{M}_{Q}$, we have
$\mathcal{M}_{Q}=\left\{\nu\in\mathcal{M}_{1}(\mathcal{G})\mid D_{\nu}\overset{\mathrm{d}}{=}D_{\mu}\right\}.$
Define $\mathcal{D}=\{Q\in\mathcal{M}_{\mathcal{B}}\mid\mathcal{M}_{Q}\cap\mathcal{R}\neq\varnothing\}$, and note that $\mathcal{R}=\bigcup_{Q\in\mathcal{D}}\mathcal{M}_{Q}$ since $\mathcal{R}$ is $\mathcal{G}$-law invariant. Define $\tau_{Q}=\tau_{\mu}$ for some $\mu\in\mathcal{M}_{Q}$ (thus all $\mu\in\mathcal{M}_{Q}$ since $(\tau_{\mu})_{\mu\in\mathcal{R}}$ is law invariant), therefore
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\}$ $=\sup_{Q\in\mathcal{D}}\sup_{\mu\in\mathcal{M}_{Q}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\}$
=(14)${=}\sup_{Q\in\mathcal{D}}\left\{\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])Q(\mathrm{d}\alpha)+\tau_{Q}(X)\right\}.$ ) } .
$(\mathrm{ii})\Rightarrow(\mathrm{i})$: The fact that $\rho$ is cash invariant, convex, and positively homogenous is clear. Let $X,Y\in L^{\infty}$ with $X\leq Y$ and $Q\in\mathcal{D}$. Then we have
$\int_{[0,1)}\mathrm{ES}_{\alpha}\left(\mathbb{E}[X|\mathcal{G}]\right)Q(\mathrm{d}\alpha)+\tau_{Q}(X)$ $=\sup_{\mu\in\mathcal{M}_{Q}}\left\{\mathbb{E}^{\mu}\left[X\right]+\tau_{Q}(X)\right\}$
$\leq\sup_{\mu\in\mathcal{M}_{Q}}\left\{\mathbb{E}^{\mu}\left[Y\right]+\tau_{Q}(Y)\right\}$
$=\int_{[0,1)}\mathrm{ES}_{\alpha}\left(\mathbb{E}[Y|\mathcal{G}]\right)Q(\mathrm{d}\alpha)+\tau_{Q}(Y).$
Therefore $\rho$ is monotonic as it is the supremum of monotonic functionals, so it is coherent.
Fix $Q\in\mathcal{D}$. The mapping from $L^{\infty}$ to $\mathbb{R}$ given by
$X\mapsto\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])Q(\mathrm{d}\alpha)$
is Fatou continuous and coherent by Propositions 11 and 15, and therefore $w(L^{\infty},\mathcal{M})$-lower semicontinuous. Thus, the functional on $L^{\infty}$ given by
$X\mapsto\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])Q(\mathrm{d}\alpha)+\tau_{Q}(X)$
is $w(L^{\infty},\mathcal{M})$-lower semicontinuous by Corollary 2, therefore $\rho$ is $w(L^{\infty},\mathcal{M})$-lower semicontinuous as it is the supremum of $w(L^{\infty},\mathcal{M})$-lower semicontinuous functionals. Since $\rho$ is coherent, it is Fatou. Let $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$ such that $\mathbb{E}[X|\mathcal{G}]\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$, let $Z=X-\mathbb{E}[X|\mathcal{G}]+Y$. For all $Q\in D$,
$\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[Z|\mathcal{G}])Q(\mathrm{d}\alpha)+\tau_{Q}(Z)$ $=\int_{[0,1)}\mathrm{ES}_{\alpha}(Y)Q(\mathrm{d}\alpha)+\tau_{Q}(X)$
$=\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])Q(\mathrm{d}\alpha)+\tau_{Q}(X),$
where the last equality follows from the fact that $\mathrm{ES}_{\alpha}$ is law invariant. Therefore, $\rho(Z)=\rho(X)$. ∎
E.5 Examples
This section will give examples of partially law-invariant and strongly partially law-invariant risk measures. We use the following lemma from Delbaen (2000, Section 4.3.i) frequently in this section.
Lemma 5.
Let $\rho_{1}:L^{\infty}\to\mathbb{R}$ and $\rho_{2}:L^{\infty}\to\mathbb{R}$ be Fatou coherent risk measures. Then $\rho_{1}\vee\rho_{2}$ is a Fatou coherent risk measure with supporting set equal to $\operatorname*{conv}(\mathcal{S}_{\rho_{1}}\cup\mathcal{S}_{\rho_{2}})$.
Let $(\Omega_{k},\mathcal{F}_{k},\mathbb{P}_{k})=([0,1],\mathcal{B}([0,1]),\lambda)$ for $k=1,2$. For the rest of the section let $(\Omega,\mathcal{F},\mathbb{P})=(\Omega_{1}\times\Omega_{2},\mathcal{F}_{1}\otimes\mathcal{F}_{2},\mathbb{P}_{1}\times\mathbb{P}_{2})$ and $\mathcal{G}=\sigma(\pi_{1})$ where $\pi_{1}(\omega,\tilde{\omega})=\omega$. Define the set $\mathcal{S}_{1}=\{\mu\in\mathcal{M}_{1}\mid D_{\mu}\in L^{1}(\mathcal{G}),D_{\mu}\leq 2\}$. This set is $w(\mathcal{M},L^{\infty})$-closed and convex and corresponds to the Fatou coherent risk measure $\rho_{1}=\mathrm{ES}_{1/2}(\mathbb{E}[X|\mathcal{G}])$.
The following example gives a Fatou $\mathcal{G}$-law-invariant coherent risk measure that is not strongly $\mathcal{G}$-law invariant.
Example 11.
Define $\nu\in\mathcal{M}_{1}$ by $D_{\nu}:\Omega\to\mathbb{R}:(\omega,\tilde{\omega})\mapsto 4\omega\tilde{\omega}$. We claim that the Fatou coherent risk measure $\rho:L^{\infty}\to\mathbb{R}:X\mapsto\mathbb{E}[D_{\nu}X]\vee\rho_{1}(X)$, which has supporting set $\mathcal{S}_{\rho}=\operatorname*{conv}(\mathcal{S}_{1}\cup\{\nu\})$, is $\mathcal{G}$-law invariant but not strongly $\mathcal{G}$-law invariant. To see this, note $\mathbb{E}[D_{\nu}|\mathcal{G}](\omega,\tilde{\omega})=2\omega$. So the measure associated to $\mathbb{E}[D_{\nu}|\mathcal{G}]$ is in the set $\mathcal{S}_{1}$. Therefore, $\operatorname*{conv}(\mathcal{S}_{1}\cup\{\nu\})^{\mathcal{G}}=\operatorname*{conv}(\mathcal{S}_{1}^{\mathcal{G}}\cup\{\nu\}^{\mathcal{G}})=\mathcal{S}_{1}$, which is $\mathcal{G}$-law invariant. Thus, $\rho$ is $\mathcal{G}$-law invariant by Theorem 1. Define the following random variables
$X(\omega,\tilde{\omega})=\omega,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ Y(\omega,\tilde{\omega})=1-\omega\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ Z(\omega,\tilde{\omega})=\tilde{\omega}-1/2.$
We have that $Z\in\ker(\mathcal{G})$, $\mathbb{E}[D_{\nu}Z]=1/6$, $X,Y\in L^{\infty}(\mathcal{G})$ and $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$, $\mathrm{ES}_{\alpha}(X)=\mathrm{ES}_{\alpha}(Y)=3/4$, $\mathbb{E}[D_{\nu}X]=2/3$ and $\mathbb{E}[D_{\nu}Y]=1/3.$ Thus, $\rho(Z+X)=5/6$ while $\rho(Z+Y)=3/4$, implying that $\rho$ is not strongly $\mathcal{G}$-law invariant by definition.
The following example gives a Fatou strongly $\mathcal{G}$-law invariant coherent risk measure that is not in the form of (7).
Example 12.
Define $\nu\in\mathcal{M}_{1}$ by $D_{\nu}:\Omega\to\mathbb{R}:(\omega,\tilde{\omega})\mapsto 2\tilde{\omega}$. We claim that the Fatou coherent risk measure $\rho:L^{\infty}\to\mathbb{R}:X\mapsto\mathbb{E}[D_{\nu}X]\vee\rho_{1}(X)$ is strongly $\mathcal{G}$-law invariant. To see this let $Z\in\ker(\mathcal{G})$ and $X,Y\in L^{\infty}(\mathcal{G})$ with $X\overset{\mathrm{d}}{=}_{\mathbb{P}}Y$. We have
$\rho(Z+X)=\rho_{1}(Z+X)\vee\mathbb{E}[D_{\nu}(Z+X)]$ $=\rho_{1}(X)\vee(\mathbb{E}[D_{\nu}Z]+\mathbb{E}[X])$
$=\rho_{1}(Y)\vee(\mathbb{E}[D_{\nu}Z]+\mathbb{E}[Y])=\rho(Z+Y).$
Note that $\widehat{\mathcal{S}}_{\rho}$ is not contained in $L^{1}(\mathcal{G})$, thus $\rho$ is not in the form of (7).
E.6 Additional results and proofs accompanying Section 5.3
To do this, we will need some results from convex analysis. Let $\left(\mathfrak{X},\mathfrak{Y},\langle\cdot,\cdot\rangle\right)$ be a dual pairing and $f:\mathfrak{X}\to\mathbb{R}\cup\{\infty\}$ such that $\operatorname*{dom}(f)=\left\{x\in\mathfrak{X}\mid f(x)<\infty\right\}\neq\varnothing$, define
$f^{*}$ $:\mathfrak{Y}\to\mathbb{R}\cup\{\infty\}:y\mapsto\sup_{x\in\mathfrak{X}}\left(\langle x,y\rangle-f(x)\right),$
$f^{**}$ $:\mathfrak{X}\to\mathbb{R}\cup\{\infty,-\infty\}:x\mapsto\sup_{y\in\mathfrak{Y}}\left(\langle x,y\rangle-f^{*}(y)\right).$
The upper epigraph of $f$ is given by $G_{f}=\left\{(x,c)\in\mathfrak{X}\times\mathbb{R}\mid f(x)\leq c\right\}\subseteq\mathfrak{X}\times\mathbb{R}$. Equip $\mathfrak{X}\times\mathbb{R}$ with the product topology $w(\mathfrak{X},\mathfrak{Y})\times E_{\mathbb{R}}$, where $E_{\mathbb{R}}$ is the standard Euclidean topology.
Proposition 22.
Let $f:\mathfrak{X}\to\mathbb{R}\cup\{\infty\}$ with $\operatorname*{dom}(f)\neq\varnothing$.
  1. (i)
    The mapping $f$ is convex if and only if $G_{f}$ is convex;
  2. (ii)
    the mapping $f$ is $w(\mathfrak{X},\mathfrak{Y})$-lower semicontinuous if and only if $G_{f}$ is $\left(w(\mathfrak{X},\mathfrak{Y})\times E_{\mathbb{R}}\right)$-closed;
  3. (iii)
[…]
  4. (iv)
    the mapping $f$ is $w(\mathfrak{X},\mathfrak{Y})$-lower semicontinuous and convex if and only if $f^{**}=f$.
See Zalinescu (2002) for a proof of Proposition 22.
First, we obtain a representation of $\mu$-convex adjustments, similar to Proposition 17.
Proposition 23.
Let $\tau:L^{\infty}\to\mathbb{R}$, $\mu\in\mathcal{M}_{1}(\mathcal{G})$, and $\rho=\mathbb{E}^{\mu}+\tau$. The following are equivalent.
  1. (i)
    The mapping $\tau$ is a Fatou $\mu$-convex adjustment;
  2. (ii)
    the mapping $\rho$ is a $w(L^{\infty},L^{1})$-lower semicontinuous convex risk measure with $\operatorname*{dom}(\rho^{*})\subseteq\mathcal{E}(\mu)$;
  3. (iii)
    $\tau$ can be represented by
    $\tau(X)=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\rho^{*}(\nu)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$ (21)
  4. (iv)
    there exists $\alpha:\mathcal{E}(\mu)\to\mathbb{R}\cup\{\infty\}$ such that $\tau$ can be represented by
    $\tau(X)=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\alpha(\nu)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Proof.
(i)$\Rightarrow$(ii): It is clear that $\rho$ is a convex risk measure and is Fatou (simple application of DCT). This implies that $\rho$ is $w(L^{\infty},L^{1})$-lower semicontinuous. Note by Föllmer and Schied (2016, Theorem 4.33), $\operatorname*{dom}(\rho^{*})\subseteq\mathcal{M}_{1}$. Let $\nu\in\mathcal{M}_{1}\backslash\mathcal{E}(\mu)$. There exists $A\in\mathcal{G}$ such that $\mathbb{E}[D_{\nu}\mathds{1}_{A}]>\mathbb{E}[D_{\mu}\mathds{1}_{A}]$. Thus, we have
$\rho^{*}(\nu)=\sup_{X\in L^{\infty}}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\tau(X)\right\}$ $\geq\sup_{\lambda\in\mathbb{R}}\left\{\lambda\mathbb{E}^{\nu}[\mathds{1}_{A}]-\lambda\mathbb{E}^{\mu}[\mathds{1}_{A}]-\tau(\lambda\mathds{1}_{A})\right\}$
$=\sup_{\lambda\in\mathbb{R}}\left\{\lambda(\mathbb{E}^{\nu}[\mathds{1}_{A}]-\mathbb{E}^{\mu}[\mathds{1}_{A}])\right\}-\tau(0)=\infty.$
Hence, $\nu\notin\operatorname*{dom}(\rho)$. (ii)$\Rightarrow$(iii): Since $\rho$ is convex and $w(L^{\infty},L^{1})$-lower semicontinuous. By Proposition 5, we have for $X\in L^{\infty}$,
$\tau(X)=\rho(X)-\mathbb{E}^{\mu}[X]=\sup_{\nu\in\operatorname*{dom}(\rho^{*})}\left\{\mathbb{E}^{\nu}[X]-\rho^{*}(\nu)\right\}-\mathbb{E}^{\mu}[X]$ $=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\rho^{*}(\nu)\right\}-\mathbb{E}^{\mu}[X]$
$=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\rho^{*}(\nu)\right\}.$
(iii)$\Rightarrow$(iv): This is trivial. (iv)$\Rightarrow$(i): The fact that $\tau$ is a $\mu$-convex adjustment is clear. Let $(X_{n})_{n=1}^{\infty}$ be a uniformly bounded sequence in $L^{\infty}$ which converges $\mathbb{P}$-a.s. to some $X\in L^{\infty}$. We have
$\tau(X)=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}[(D_{\nu}-D_{\mu})X]-\alpha(\nu)\right\}$ $=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\liminf\limits_{n\to\infty}\left(\mathbb{E}[(D_{\nu}-D_{\mu})X_{n}]-\alpha(\nu)\right)\right\}$
$\leq\liminf\limits_{n\to\infty}\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}[(D_{\nu}-D_{\mu})X]-\alpha(\nu)\right\}$
$=\liminf\limits_{n\to\infty}\tau(X_{n}).$
Thus proving the Fatou continuity of $\tau$. ∎
For $f:\mathfrak{X}\to\mathbb{R}\cup\{\infty\}$ let $L_{f}^{\leq c}=\left\{x\in\mathfrak{X}\mid f(x)\leq c\right\}$, $L_{f}^{<c}=\left\{x\in\mathfrak{X}\mid f(x)<c\right\}$, and $L_{f}^{=c}=\left\{x\in\mathfrak{X}\mid f(x)=c\right\}$. We will also need the following two lemmas.
Lemma 6.
Let $\left(\mathfrak{X},\mathfrak{Y},\langle\cdot,\cdot\rangle\right)$ be a dual pairing and $f:\mathfrak{X}\to\mathbb{R}\cup\{\infty\}$ be convex with $\operatorname*{dom}(\rho)\neq\varnothing$. Then $\overline{L^{\leq c}_{f}}=L^{\leq c}_{f^{**}}$ for all $c>-f^{*}(0)=\inf_{x\in\mathfrak{X}}f(x)=\inf_{x\in\mathfrak{X}}f^{**}(x)$.
Proof.
Let $c>-f^{*}(0)$, since $f^{**}(x)\leq f(x)$ for all $x\in\mathfrak{X}$, $L^{\leq c}_{f}\subseteq L^{\leq c}_{f^{**}}$. We have $\overline{L^{\leq c}_{f}}\subseteq L^{\leq c}_{f^{**}}$ since $f^{**}$ is $w(\mathfrak{X},\mathfrak{Y})$-lower semicontinous as it is the supremum of $w(\mathfrak{X},\mathfrak{Y})$-continuous functions (meaning $L^{\leq c}_{f^{**}}$ is $w(\mathfrak{X},\mathfrak{Y})$-closed).
Let $\epsilon>0$, we claim that $L^{\leq c-\epsilon}_{f^{**}}\subseteq\overline{L^{\leq c}_{f}}$. Clearly this is true if $L^{\leq c-\epsilon}_{f^{**}}=\varnothing$. Therefore assume that $L^{\leq c-\epsilon}_{f^{**}}\neq\varnothing$ and let $x\in L^{\leq c-\epsilon}_{f^{**}}$, which means $(x,c-\epsilon)\in G_{f^{**}}$. There exists $(x_{\lambda},c_{\lambda})\in G_{f}$ such that $\lim_{\lambda}(x_{\lambda},c_{\lambda})=(x,c-\epsilon)$ since $\overline{G_{f}}=G_{f^{**}}$. Find $\lambda_{0}$ such that for all $\lambda\geq\lambda_{0}$, $c_{\lambda}\leq c$. Therefore for all $\lambda\geq\lambda_{0}$ we have $x_{\lambda}\in L_{f}^{\leq c}$. Thus the tail net $(x_{\lambda})_{\lambda\geq\lambda_{0}}$ is contained in $L_{f}^{\leq c}$ and converges to $x$, therefore $x\in\overline{L_{f}^{\leq c}}$. This shows the claim that $L^{\leq c-\epsilon}_{f^{**}}\subseteq\overline{L_{f}^{\leq c}}$. Since $L^{\leq c-\epsilon}_{f^{**}}\subseteq\overline{L_{f}^{\leq c}}$ for arbitrary $\epsilon>0$, we have $L^{<c}_{f^{**}}\subseteq\overline{L_{f}^{\leq c}}$.
Clearly $L^{<c}_{f^{**}}\subseteq L^{\leq c}_{f^{**}}$, implying $\overline{L^{<c}_{f^{**}}}\subseteq L^{\leq c}_{f^{**}}$. Let $x\in L^{\leq c}_{f^{**}}$ and $y\in L^{<c}_{f^{**}}$ (non-empty since $c>-f^{*}(0)$), define
$z_{n}=\left(1-\frac{1}{n}\right)x+\frac{1}{n}y,$
which is clearly in $L^{<c}_{f^{**}}$ by the convexity of $f^{**}$. Since $\lim_{n}z_{n}=x$, $x\in\overline{L^{<c}_{f^{**}}}$. Showing that $\overline{L^{<c}_{f^{**}}}=L^{\leq c}_{f^{**}}$. Therefore
$L^{\leq c}_{f^{**}}=\overline{L^{<c}_{f^{**}}}\subseteq\overline{L_{f}^{\leq c}}\subseteq L^{\leq c}_{f^{**}}.$
Which completes the proof. ∎
Lemma 7.
Let $\rho:L^{\infty}\to\mathbb{R}$ be a Fatou convex risk measure. Then $\rho$ is law invariant if and only if $L^{\leq c}_{\rho^{*}}$ is law invariant for all $c>-\rho(0)$.
Proof.
By Theorem 4.59 in Föllmer and Schied (2016), $\rho$ is law invariant if and only if $\rho^{*}$ is law invariant. Clearly, if $\rho^{*}$ is law invariant, $L^{\leq c}_{\rho^{*}}$ is law invariant for all $c>-\rho(0)$. Now assume that $L^{\leq c}_{\rho^{*}}$ is law invariant for all $c>-\rho(0)$. Fix some $c>-\rho(0)=\inf_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\rho^{*}(\mu)$, find $N\in\mathbb{N}$ such that $c-\frac{1}{N}>-\rho(0)$. We have
$L^{=c}_{\rho^{*}}=\cap_{n\geq N}\left(L^{\leq c}_{\rho^{*}}/L^{\leq c-\frac{1}{n}}_{\rho^{*}}\right)$
is law invariant. Now let $c=-\rho(0)$, we have
$L^{=c}_{\rho^{*}}=\cap_{n\geq 1}\left(L^{\leq c+\frac{1}{n}}_{\rho^{*}}\right)$
is law invariant (since for $c<-\rho(0)$, $L^{\leq c}_{\rho^{*}}=\varnothing$). This shows that $\rho^{*}$ is law invariant. ∎
Proof of Theorem 5.
Let $\rho$ be a Fatou convex risk measure. By Föllmer and Schied (2016, Theorem 4.18, Theorem 4.33), we have
$\rho(X)=\sup_{\nu\in\mathcal{M}_{1}}\left\{\mathbb{E}^{\nu}[X]-\rho^{*}(\nu)\right\}$ $=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\rho^{*}(\nu)\right\}$
$=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left\{\mathbb{E}^{\mu}[X]+\sup_{\nu\in\mathcal{E}(\mu)}\left(\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\rho^{*}(\nu)\right)\right\}$
$=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{E}(\mu)}\left\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]-\rho^{*}(\nu)\right\}$, $X\in L^{\infty}$. Note that for all $\mu\in\mathcal{M}_{1}(\mathcal{G})$, $\tau_{\mu}=-\infty$ on $L^{\infty}$ or $\operatorname*{dom}(\tau_{\mu})=L^{\infty}$. By Proposition 23, $\tau_{\mu}$ is a Fatou $\mu$-convex adjustment. Define $N:\mathcal{M}_{1}(\mathcal{G})\to\mathbb{R}\cup\{\infty\}:\mu\mapsto-\tau_{\mu}(0)$. We claim that $N$ is convex, let $\mu_{1},\mu_{2}\in\mathcal{M}_{1}(\mathcal{G})$ and $\lambda\in[0,1]$. Define $\upsilon=\lambda\mu_{1}+(1-\lambda)\mu_{2}$, we have
$N\left(\upsilon\right)=\inf_{\nu\in\mathcal{E}(\upsilon)}\rho^{*}\left(\nu\right)$ $\leq\inf_{(\nu_{1},\nu_{2})\in\mathcal{E}(\mu_{1})\times\mathcal{E}(\mu_{2})}\rho^{*}\left(\lambda\nu_{1}+(1-\lambda)\nu_{2}\right)$
$\leq\lambda\inf_{\nu_{1}\in\mathcal{E}(\mu_{1})}\rho^{*}\left(\nu_{1}\right)+(1-\lambda)\inf_{\nu_{2}\in\mathcal{E}(\mu_{2})}\rho^{*}\left(\nu_{2}\right)$
$=\lambda N(\mu_{1})+(1-\lambda)N(\mu_{2}),$
where we are using the convexity of $\rho^{*}$. This completes the forward direction. The reverse direction is a simple application of Proposition 23, where one expands the convex adjustments by (21). To see the last claim, define $\tilde{\rho}=\rho|_{L^{\infty}(\mathcal{G})}$, note
$\tilde{\rho}(X)=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left(\mathbb{E}^{\mu}[X]-N(\mu)\right)=N^{*}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G}).$
Thus $N^{**}=\tilde{\rho}^{*}$ and by Lemma 6, for all $c>-\rho(0)$, $\overline{L^{\leq c}_{N}}=L^{\leq c}_{\tilde{\rho}^{*}}$. Using Lemma 7,
$\rho\text{ is }\mathcal{G}\text{-law invariant}\Leftrightarrow\tilde{\rho}\text{ is law invariant}$ $\Leftrightarrow L^{\leq c}_{\tilde{\rho}^{*}}\text{ is }\mathcal{G}\text{-law invariant for all }c>-\rho(0)$
$\Leftrightarrow\overline{L^{\leq c}_{N}}\text{ is }\mathcal{G}\text{-law invariant for all }c>-\rho(0)$
$\Leftrightarrow L^{\leq c}_{N}\text{ is almost }\mathcal{G}\text{-law invariant for all }c>-\rho(0).$
Thus proving the claim. ∎
Appendix F Additional results and proofs accompanying Section 6
We call $\mathcal{X}$ upward-directed if for all $X_{1},X_{2}\in\mathcal{X}$, $\max\{X_{1},X_{2}\}\in\mathcal{X}$.
Proposition 24 (Föllmer and Schied (2016, Theorem A.11)).
Let $\mathcal{X}\subseteq L^{\infty}$ be such that $\sup_{X\in\mathcal{X}}\|X\|_{\infty}$ is finite. Then $\operatorname*{ess\text{-}sup}\mathcal{X}$ exists. Moreover, if $\mathcal{X}$ is upward-directed, then it holds that $\mathbb{E}\left[\operatorname*{ess\text{-}sup}\mathcal{X}\right]=\sup_{X\in\mathcal{X}}\mathbb{E}[X]$.
Lemma 8.
Let $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$ be $\mathcal{G}$-convex. For all $X\in L^{\infty}$,
$\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X]=\mathbb{E}\left[\operatorname*{ess\text{-}sup}_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X|\mathcal{G}]\right].$
Proof.
Let $X\in L^{\infty}$, define $I(X)=\{\mathbb{E}^{\mu}[X|\mathcal{G}]:\mu\in\mathcal{Q}\}$. We claim that $I(X)$ is upward-directed. Given $\mu_{1},\mu_{2}\in\mathcal{Q}$, let
$A=\left\{\omega:\mathbb{E}^{\mu_{1}}[X|\mathcal{G}](\omega)\geq\mathbb{E}^{\mu_{2}}[X|\mathcal{G}](\omega)\right\}.$
Since $\{D_{\mu}:\mu\in\mathcal{Q}\}$ is $\mathcal{G}$-convex, $\mathds{1}_{A}D_{\mu_{1}}+\mathds{1}_{A^{c}}D_{\mu_{2}}\in\{D_{\mu}:\mu\in\mathcal{Q}\}$. Let $\nu$ be the probability measure associated with $\mathds{1}_{A}D_{\mu_{1}}+\mathds{1}_{A^{c}}D_{\mu_{2}}$. We have
$\mathbb{E}^{\nu}[X|\mathcal{G}]=I_{A}\mathbb{E}^{\mu_{1}}[X|\mathcal{G}]+I_{A^{c}}\mathbb{E}^{\mu_{2}}[X|\mathcal{G}]=\max\left\{\mathbb{E}^{\mu_{1}}[X|\mathcal{G}],\mathbb{E}^{\mu_{2}}[X|\mathcal{G}]\right\}.$
Since $\mathbb{E}^{\nu}[X|\mathcal{G}]\in I(X)$, $I(X)$ is upward directed. By Proposition 24,
$\mathbb{E}[\operatorname*{ess\text{-}sup}I(X)]=\sup_{\mu\in\mathcal{Q}}\mathbb{E}[\mathbb{E}^{\mu}[X|\mathcal{G}]]=\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[\mathbb{E}^{\mu}[X|\mathcal{G}]]=\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X].$
Since $X\in L^{\infty}$ was arbitrary, the proof is completed. ∎
Proof of Proposition 9.
Fix some $X\in L^{\infty}$, we have
$\overline{\mathrm{ES}}_{\alpha}^{\mathcal{Q}}[X]=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}[X]=\sup_{\mu\in\mathcal{Q}}\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}\left[(X-x)^{+}\right]\right).$ (22)
Define $f:\mathcal{Q}\times\mathbb{R}\to\mathbb{R}:(\mu,x)\mapsto x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}[(X-x)^{+}]$. Since $\mathcal{Q}$ is $\mathcal{G}$-convex, it is convex. Clearly, $f(\cdot,x)$ is $w(\mathcal{M},L^{\infty})|_{\mathcal{Q}}$-continuous and affine for all $x\in\mathbb{R}$. Likewise, $f(\mu,\cdot)$ is continuous and convex for all $\mu\in\mathcal{Q}$. By the minimax Theorem from Mertens et al. (2015, Theorem I.1.1), we are able to switch the minimum and supremum in (22), thus
$\overline{\mathrm{ES}}_{\alpha}^{\mathcal{Q}}[X]=\min_{x\in\mathbb{R}}\sup_{\mu\in\mathcal{Q}}\left(x+\frac{1}{1-\alpha}\mathbb{E}^{\mu}\left[(X-x)^{+}\right]\right)$ $=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\sup_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}\left[(X-x)^{+}\right]\right)$
$=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}\left[\operatorname*{ess\text{-}sup}_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}\left[(X-x)^{+}|\mathcal{G}\right]\right]\right)$
$=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}\left[\rho_{\mathcal{G}}^{\mathcal{Q}}\left((X-x)^{+}\right)\right]\right),$
where the third equality follows from Lemma 8.
Let $\mu_{1},\mu_{2}\in\mathcal{Q}_{\beta}$ and $\lambda\in L^{\infty}(\mathcal{G})$ with $0\leq\lambda\leq 1$, clearly
$0\leq\lambda D_{\mu_{1}}+(1-\lambda)D_{\mu_{2}}\leq\frac{1}{1-\beta}.$
Let $A\in\mathcal{G}$,
$\mathbb{E}[(\lambda D_{\mu_{1}}+(1-\lambda)D_{\mu_{2}})\mathds{1}_{A}]$ $=\mathbb{E}[\lambda\mathds{1}_{A}\mathbb{E}[D_{\mu_{1}}|\mathcal{G}]+(1-\lambda)\mathds{1}_{A}\mathbb{E}[D_{\mu_{1}}|\mathcal{G}]]$
$=\mathbb{E}[\lambda\mathds{1}_{A}+(1-\lambda)\mathds{1}_{A}]=\mathbb{P}(A).$
Therefore, $\lambda D_{\mu_{1}}+(1-\lambda)D_{\mu_{2}}\in\widehat{\mathcal{Q}}_{\beta}$. ∎
A computable formula for (16).
Let $\Phi$ (resp. $\phi$) denote the standard normal cdf (resp. pdf), for $X\sim N(r,\sigma^{2})$, we have
$\mathrm{VaR}_{\beta}\left((X-x)^{+}\right)$ $=\left(r-x+\sigma\Phi^{-1}(\beta)\right)^{+},$
$\mathrm{ES}_{\beta}\left((X-x)^{+}\right)$ =$=\begin{cases}r-x+\sigma\frac{\phi(\Phi^{-1}(\beta))}{1-\beta}&x\leq r+\sigma\Phi^{-1}(\beta)\\ \frac{(r-x)(1-\Phi\left(\frac{x-r}{\sigma}\right))}{1-\beta}+\sigma\frac{\phi(\frac{x-r}{\sigma})}{1-\beta}&x>r+\sigma\Phi^{-1}(\beta).\end{cases}$
For $l\in\mathbb{R}$, the conditional distribution
$\pi_{1}X_{1}+\pi_{2}X_{2}|X_{1}=l\sim N\left(\pi_{1}l+\pi_{2}m_{2}+c\pi_{2}\frac{\sigma_{2}}{\sigma_{1}}(l-m_{1}),\pi_{2}^{2}(1-c^{2})\sigma_{2}^{2}\right).$
Define the functions $M(\pi_{1},\pi_{2},l)=\pi_{1}l+\pi_{2}m_{2}+c\pi_{2}\frac{\sigma_{2}}{\sigma_{1}}(l-m_{1})$, $\sigma(\pi_{2})=\pi_{2}\sqrt{(1-c^{2})}\sigma_{2}$, and
$g_{\beta}(r,\sigma,x)=\begin{cases}r-x+\sigma\frac{\phi(\Phi^{-1}(\beta))}{1-\beta}&x\leq r+\sigma\Phi^{-1}(\beta)\\ \frac{(r-x)(1-\Phi\left(\frac{x-r}{\sigma}\right))}{1-\beta}{}+\sigma\frac{\phi(\frac{x-r}{\sigma})}{1-\beta}&x>r+\sigma\Phi^{-1}(\beta).\end{cases}$
We can now verify the following identity
$\mathrm{ES}_{\beta}((\pi_{1}X_{1}+\pi_{2}X_{2}-x)^{+}|X_{1})=g_{\beta}(M(\pi_{1},\pi_{2},X_{1}),\sigma(\pi_{2}),x).$ (23)
Therefore,
$f_{\beta}(\pi_{1},\pi_{2},x)=\int_{\mathbb{R}}g_{\beta}(M(\pi_{1},\pi_{2},z),\sigma(\pi_{2}),x)\leavevmode\nobreak\ \frac{1}{\sigma_{1}}\phi\left(\frac{z-m_{1}}{\sigma_{1}}\right)\leavevmode\nobreak\ \mathrm{d}z,$
which is an integral formula for (16) that is easy to numerically compute. ∎
Appendix G Analysis in a finite probability space
Will first show that law invariance is equivalent to permutation invariance in a finite uniform probability space. Afterward, we will study partially law-invariant coherent risk measures when our probability space has a natural structure.
G.1 Law-invariant coherent risk measures
In the literature, when proving results regarding law-invariant risk measures, one often assumes that the underlying probability space is atomless. Since finite probability spaces always have atoms, we instead will assume that the underlying probability measure is uniform.
Let $N\in\mathbb{N}$, denote by $[N]=\{1,\dots,N\}$. Assume that $\Omega=[N]$ and $\mathcal{F}=2^{\Omega}$ and put the uniform probability measure $\mathbb{P}$ on $(\Omega,\mathcal{F})$, that is ($\mathbb{P}(\{k\})=1/N$ for all $k\in[N]$. We can identify elements $X\in L^{\infty}$ with $\mathbb{R}^{N}$. For the rest of the section, we will refer only to $\mathbb{R}^{N}$, but this is to be regarded as $L^{\infty}$.
Let $\Delta^{N}=\left\{\mu\in\mathbb{R}_{+}^{N}\hskip 2.0pt:\hskip 2.0pt\mathbb{E}[\mu]=1\right\}$, it is easy to see that we can identify $\mathcal{M}_{1}$ with $\Delta^{N}$. Since we are working in $\mathbb{R}^{N}$, the $w\left(\mathbb{R}^{N},\mathbb{R}^{N}\right)$ topology ($w\left(L^{\infty},\mathcal{M}\right)$ and $w\left(\mathcal{M},L^{\infty}\right)$) is the standard euclidean topology. Therefore, there will be no mention in this section as to which topology we are dealing with. Note that any closed set $\mathcal{R}\subseteq\Delta^{N}$ is compact. Finally, Fatou continuity is implied by any coherent risk measure $\rho:\mathbb{R}^{N}\to\mathbb{R}$. Thus, there will be no mention of continuity in this section.
To describe law invariance in this space define $S_{N}=\{\sigma:[N]\to[N]\mid\sigma\text{ is a bijection}\}$. For $\sigma\in S_{N}$ define the map $T_{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$ where $(T_{\sigma}(X))_{n}=X_{\sigma(n)}$.
Definition 4.
We say that a mapping $\rho:\mathbb{R}^{N}\to\mathbb{R}$ is permutation invariant if $\rho(T_{\sigma}(X))=\rho(X)$ for all $\sigma\in S_{N}$ and $X\in\mathbb{R}^{N}$. We say a set $\mathcal{R}\subseteq\mathbb{R}^{N}$ is permutation invariant if $T_{\sigma}(\mathcal{R})\subseteq\mathcal{R}$ for all $\sigma\in S_{N}$.
It is easy to see that a mapping is law invariant if and only if it is permutation invariant. If $\mathcal{R}\subseteq\mathbb{R}^{N}$ is permutation invariant then $T_{\sigma}(\mathcal{R})\supseteq T_{\sigma}(T_{\sigma^{-1}}(\mathcal{R}))=\mathcal{R}$ so we actually have equality.
Proposition 25.
Let $\rho:\mathbb{R}^{N}\to\mathbb{R}$ be a coherent risk measure. Then $\rho$ is law invariant if and only if $\mathcal{S}_{\rho}$ is permutation invariant.
Proof.
Let $\mathcal{S}_{\rho}$ be permutation invariant then we have
$\rho(T_{\sigma}(X))=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}[\mu T_{\sigma}(X)]$ $=\frac{1}{N}\sup_{\mu\in\mathcal{S}_{\rho}}(\mu\cdot T_{\sigma}(X))$
$=\frac{1}{N}\sup_{\mu\in\mathcal{S}_{\rho}}(T_{\sigma^{-1}}(\mu)\cdot X)=\frac{1}{N}\sup_{\mu\in\mathcal{S}_{\rho}}(\mu\cdot X)=\rho(X),$
for all $\sigma\in S_{n}$. Conversely, assume that $\mathcal{S}_{\rho}$ is not permutation invariant. Find $R\in\mathcal{S}_{\rho}$ and $\sigma\in S_{n}$ such that $T_{\sigma}(R)\notin\mathcal{S}_{\rho}$. Since $\mathcal{S}_{\rho}$ is compact and convex, by the Hahn Banach separation theorem, we can find $X\in\mathbb{R}^{N}$ such that
$\rho(X)=\frac{1}{N}\sup_{\mu\in\mathcal{S}_{\rho}}(\mu\cdot X)<\frac{1}{N}X\cdot T_{\sigma}(R).$
Therefore we have
$\rho(X)<\frac{1}{N}X\cdot T_{\sigma}(R)=\frac{1}{N}T_{\sigma^{-1}}(X)\cdot R\leq\rho(T_{\sigma^{-1}}(X)).$
Implying that $\rho$ is not permutation invariant and thus not law invariant. ∎
G.2 Partially law-invariant coherent risk measures
We write $A\times B=\{(a,b):a\in A,\leavevmode\nobreak\ b\in B\}$ for two sets $A$ and $B$.
Let $\Omega=[M]\times[N]$ for some $M,N\in\mathbb{N}$, $\mathcal{F}=2^{\Omega}$, $\mathcal{G}$ is generated by the projection map onto the second coordinate, and $\mathbb{P}$ is positive (that is, $\mathbb{P}>0$ on ${}\mathcal{F}\setminus\{\varnothing\}$) and uniform in the second coordinate (that is, $\mathbb{P}([M]\times\{n\})=1/N$ for all $n\in[N]$).
We can identify elements $X\in L^{\infty}$ with $\mathbb{R}^{M\times N}$. For the rest of the section, we will refer only to $\mathbb{R}^{M\times N}$ but this is to be regarded as $L^{\infty}$. Let $\Delta^{M\times N}=\{\mu\in\mathbb{R}^{M\times N}_{+}\hskip 2.0pt:\hskip 2.0pt\mathbb{E}[\mu]=1\}$, it is easy to see that $\mathcal{M}_{1}$ can be identified with $\Delta^{M\times N}$. The same comments about topology and Fatou continuity from the previous section also apply in this section.
For $k\in[N]$ define $E^{k}\in\mathbb{R}^{M\times N}$ by $E^{k}_{mn}=\delta_{kn}$, where $\delta_{kn}$ is the Kronecker delta. Let =$\mathcal{E}=\left\{E^{k}\hskip 2.0pt:\hskip 2.0ptk\in[N]\right\}$, clearly $\mathcal{E}$ is a basis for $L^{\infty}(\mathcal{G})$. Define the map $\mathbb{L}:\mathbb{R}^{M\times N}\to\mathbb{R}^{N}:X\mapsto[\mathbb{E}[X|\mathcal{G}]]_{\mathcal{E}}$. Given some $X\in\mathbb{R}^{N}$ we will define $\tilde{X}\in\mathbb{R}^{M\times N}$ by $\tilde{X}=\sum_{k=1}^{N}X_{k}E^{k}$. A set $\mathcal{R}\subseteq\mathbb{R}^{M\times N}$ is $\mathcal{G}$-law invariant if $\mathbb{L}(\mathcal{R})$ is permutation invariant.
The following proposition gives necessary and sufficient conditions for the supporting set of an $\mathcal{G}$-law invariant coherent risk measure.
Proposition 26.
Let $\rho:\mathbb{R}^{M\times N}\to\mathbb{R}$ be a coherent risk measure. Then $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{S}_{\rho}$ is $\mathcal{G}$-law invariant.
Proof.
Define the mapping
$\tilde{\rho}:\mathbb{R}^{N}\to\mathbb{R}:X\mapsto\rho(\tilde{X}).$
It is easy to verify that this mapping is a coherent risk measure. Let $\mathcal{S}_{\tilde{\rho}}\subseteq\Delta^{N}$ be the supporting set for $\tilde{\rho}$. It follows that $\rho$ is $\mathcal{G}$-law invariant if and only if $\tilde{\rho}$ is law invariant. For $X\in\mathbb{R}^{N}$ we have that
$\tilde{\rho}(X)=\rho(\tilde{X})=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}\left[\mu\tilde{X}\right]$ $=\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{E}\left[\mathbb{E}[\mu|\mathcal{G}]\tilde{X}\right]$
$=\frac{1}{N}\sup_{\mu\in\mathcal{S}_{\rho}}\mathbb{L}(\mu)\cdot X=\frac{1}{N}\sup_{\mu\in\mathbb{L}(\mathcal{S}_{\rho})}\mu\cdot X.$
Thus $\mathbb{L}(\mathcal{S}_{\rho})\subseteq\mathcal{S}_{\tilde{\rho}}$ by Proposition 10. Since $\mathbb{L}(\mathcal{S}_{\rho})$ is compact and convex (image of a compact and convex set under a continuous linear map), the same theorem states that $\mathbb{L}(\mathcal{S}_{\rho})=\mathcal{S}_{\tilde{\rho}}$. We then have
$\rho\text{ is }\mathcal{G}\text{-law invariant }\Leftrightarrow\tilde{\rho}\text{ is law invariant}$ $\Leftrightarrow\mathcal{S}_{\tilde{\rho}}\text{ is permutation invariant}$
$\Leftrightarrow\mathbb{L}(\mathcal{S}_{\rho})\text{ is permutation invariant}$
[…]
Thus proving that $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{S}_{\rho}$ is $\mathcal{G}$-law invariant. ∎
G.3 Representation results
This section provides representation results similar to Theorem 1 and Theorem 3. A natural corollary to Proposition 26 is the following Theorem.
Theorem 6.
The mapping $\rho:\mathbb{R}^{M\times N}\to\mathbb{R}$ is a $\mathcal{G}$-law-invariant coherent risk measure if and only if
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}[\mu X],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}$
holds for some closed and convex $\mathcal{S}\subseteq\mathcal{M}_{1}$ such that $\mathbb{L}(\mathcal{S})$ is $\mathcal{G}$-law invariant.
Define the set $\Delta_{\mathcal{G}}^{M\times N}=\Delta^{M\times N}\cap L^{\infty}(\mathcal{G})$, which can be identified with $\mathcal{M}_{1}(\mathcal{G})$.
Theorem 7.
The mapping $\rho:\mathbb{R}^{M\times N}\to\mathbb{R}$ is a coherent risk measure if and only if there exists a unique compact and convex set $\mathcal{R}\subseteq\Delta_{\mathcal{G}}^{M\times N}$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ such that
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left(\mathbb{E}\left[\mu X\right]+\tau_{\mu}(X)\right),\hskip 3.0ptX\in L^{\infty}.$ (24)
Moreover, in (24), $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{R}$ is $\mathcal{G}$-law invariant.
Proof.
Let $\rho$ be a coherent risk measure. Define the set =$\mathcal{R}=\left\{\mathbb{E}[\mu|\mathcal{G}]:\mu\in\mathcal{S}_{\rho}\right\}\subseteq\mathcal{R}\subseteq\Delta_{\mathcal{G}}^{M\times N}$. The set $\mathcal{R}$ is compact and convex since it is the image of a compact convex set under a (continuous) linear map. For $\mu\in\mathcal{R}$, define $\mathcal{S}(\mu)=\left\{\nu\in\mathcal{S}_{\rho}:\mathbb{E}[\nu|\mathcal{G}]=\mu\right\}$. It is clear that $\mathcal{S}_{\rho}=\bigcup_{\mu\in\mathcal{R}}\mathcal{S}(\mu)$. We have
$\rho(X)=\sup_{\nu\in\mathcal{S}_{\rho}}\mathbb{E}[\mu X]=\sup_{\mu\in\mathcal{R}}\sup_{\nu\in\mathcal{S}(\mu)}\mathbb{E}[\nu X]$ $=\sup_{\mu\in\mathcal{R}}\sup_{\nu\in\mathcal{S}(\mu)}(\mathbb{E}[\mu X]+\mathbb{E}[(\nu-\mu)X])$
$=\sup_{\mu\in\mathcal{R}}\left(\mathbb{E}[\mu X]+\tau_{\mu}(X)\right)$
for all $X\in L^{\infty}$, where $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\mathbb{E}[(\nu-\mu)X]$. The fact that $\tau_{\mu}\in\mathrm{CA}(\mu)$ follows a similar argument as the proof of Proposition 17. Since $\mathcal{R}$ supports $\rho|_{L^{\infty}(\mathcal{G})}$ and is closed and convex, it is unique from Proposition 10. For the converse, note that (12) is the supremum of coherent risk measures (by Proposition 8; see (11)) and thus a coherent risk measure.
To prove the last claim, let $\tilde{\rho}:\mathbb{R}^{N}\to\mathbb{R}:X\mapsto\rho(\tilde{X})$ and $\mathcal{S}_{\tilde{\rho}}\subseteq\Delta^{N}$. We have
$\tilde{\rho}(X)=\rho(\tilde{X})=\sup_{\mu\in\mathcal{R}}\mathbb{E}[\mu\tilde{X}]=\frac{1}{N}\sup_{\mu\in\mathbb{L}(\mathcal{R})}\mu\cdot X$
for all $X\in\mathbb{R}^{N}$. Therefore $\mathbb{L}(\mathcal{R})=\mathcal{S}_{\tilde{\rho}}$, and
[…]
Thus proving that $\rho$ is $\mathcal{G}$-law invariant if and only if $\mathcal{R}$ is $\mathcal{G}$-law invariant. ∎